### v1

# The Failed Migration of Academic Twitter

Abstract
Following change in Twitter’s ownership and subsequent changes to content moderation policies, many in academia looked to move their discourse elsewhere and migration to Mastodon was pursued by some. Our study looks at the dynamics of this migration. Utilizing publicly available user account data, we track the posting activity of academics on Mastodon over a one year period. Our analyses reveal significant challenges sustaining user engagement on Mastodon due to its decentralized structure as well as competition from other platforms such as Bluesky and Threads. The movement lost momentum after an initial surge of enthusiasm as most users did not maintain their activity levels, and those who did faced lower levels of engagement compared to Twitter. Our findings highlight the challenges involved in transitioning professional communities to decentralized platforms, emphasizing the need for focusing on migrating social connections for long-term user engagement.
Keywords:
Social Media Mastodon Twitter Social Movement Computational Social Science Decentralized Platforms.
**footnotetext: These authors contributed equally to this work.
1 Introduction
Social media platforms have become essential outlets for communication, information dissemination, and professional networking. Among these platforms, Twitter (now known as “X”) has gained significant popularity, often referred to as virtual town square for its role facilitating public discourse [10]. Academics and journalists populate sub-communities on Twitter and use the platform for professional communication [7, 9]. Academics have extensively used Twitter to share research findings, announce academic events such as workshops and conferences, and contribute to wider scholarly discourse [11, 8]. However, after Elon Musk’s acquisition of the company and controversial changes to content moderation policies, many on so-called Academic Twitter sought alternatives [2].
Among academics, Mastodon, a decentralized social media platform, emerged as a popular alternative with some creating servers themed around their topics of interest. Unlike traditional platforms, Mastodon operates on a federated model where independent servers interact through a common protocol, allowing users to join different instances based on their preferences while still communicating across the entire network [14]. This structure aims to reduce centralized control over internet services by distributing data and services across multiple nodes, enhancing privacy, security, and resilience against censorship and corporate control [12]. Yet despite these advantages, a large-scale measurement study of Mastodon revealed that this decentralized structure also leads to lower user engagement compared to centralized platforms like Twitter, where network effects and a larger, more active user base drive higher engagement levels [12].
The migration of users between social media platforms has been a topic of interest among researchers, particularly in the context of changes in platform policies, user experiences, and factors that influence user decisions and the implications for online communities. Social media movements often occur in response to significant changes in platform policies, ownership, user dissatisfaction or perceived deficiencies in a platform[13]. Qualitative analyses, including interviews and surveys of users migrating from platforms like Fandom to alternatives such as Tumblr, reveal that policy and value-based aspects are primary catalysts in the migration process [3].
Additionally, research has shown that user migration is influenced by factors such as perceived platform attractiveness and peer influence [5]. Study of users migrating from Twitter to Mastodon post-Elon Musk’s acquisition indicated that social influence, network structure, commitment, and shared identity significantly drive collective behavior changes in online communities [2]. Analysis showed Twitter-to-Mastodon migration was characterized by low Twitter account deletion rates, concentration of users in large Mastodon instances, varied posting behaviors across platforms [4]. A similar study on Twitter-to-Mastodon migration revealed lower user engagement and posting activity on Mastodon, highlighting challenges in retaining active users due to its decentralized architecture [6].
Despite prior work, there remains a gap in the research concerning specific professional communities, as previous work has focused on the general population of social media platform users. In this study, we address this gap by examining the migration patterns and behaviors of academic users, providing a more nuanced understanding of how these particular communities navigate platform transitions. Our study addresses the following research questions:
RQ1: To what extent were academics engaged in the migration from Twitter to Mastodon, and how can the success or failure of this migration be assessed?
RQ2: Which academic disciplines have demonstrated the highest levels of persistence on Mastodon, and what factors characterize their migration experiences?
Our findings indicate that majority of the academic researchers moved to Mastodon failed to maintain their posting activity and follower engagement. These results underscore the importance of not only transferring accounts but also migrating social connections to ensure the success of social media transitions.
2 Data
We utilized a public Github repository containing lists of academics on Mastodon [^1] https://github.com/nathanlesage/academics-on-mastodon?tab=readme-ov-file. These lists were curated by researchers from 50 distinct disciplines, who volunteered to be added to the lists. Each list corresponds to a specific research discipline, containing account details of researchers within that research area. If a list for a particular discipline does not exist, one can start a new list and provide a corresponding form. This form allows researchers to add themselves to the relevant list based on the discipline they identify with. While the database continues to expand, for the purpose of tracking, we focused only on scholars whose Mastodon accounts were already in the repository at the beginning of November 2022, noting the Twitter-to-Mastodon migration began in late October 2022. In total, we tracked the Mastodon profiles and public metrics of 7542 academic users using a customized Mastodon API. Data was collected weekly from November 2022 to October 2023 allowing for temporal analysis of the user activity and engagement over the span of a year.
For cross-platform analysis, we also collected tweets via Zeeschuimer [^2] https://github.com/digitalmethodsinitiative/zeeschuimer using the query "mastodon (researcher OR researchers OR academic OR academics OR professor OR professors OR scholar OR scholars) lang:en until:2023-11-01 since:2022-10-01". Since many tweets were included because the username contained the word "mastodon" (as some users use their Mastodon account name as their Twitter username, we further filtered out these tweets). We retained only those that contained relevant keywords in the main body of the tweet, resulting in a total of 1,497 tweets.
3 Results
3.1 Attrition of academic accounts on Mastodon
Figure 1 displays a count heat map illustrating the relationship between “last status time" and “record time”, aggregated by month. The “last status time” refers to the time of the user’s most recent post, while “record time" indicates the time at which the data was collected.
In the initial month of data collection (November 2022), there were 7,505 active scholars who had posted within that same month. This number experienced a gradual decline over the subsequent seven months, showing a minor resurgence in July 2023 before continuing its descent, ultimately reaching 2,398 by October 2023. The figure also shows that 1,237 scholars were active in November 2022 but then not again throughout the remainder of our data collection. We observe a significant rate of attrition; each month, approximately 10% to 20% of scholars in our set stopped actively using the platform.
Figure 1: Heat map of user activity. Last status vs. record time (aggregated by month).
3.2 Engagement with academic accounts on Mastodon
Given the central role of engagement in user activity, we look at the engagement with migrated academic accounts on Mastodon. To do, so we first characterize accounts based on their activity on the platform.
3.2.1 Active users.
We define active users as scholars who engaged on Mastodon during the initial Twitter-to-Mastodon migration, which began in late October 2022. For our purposes, we’ve defined this initial period between October 15 and November 30, 2022. We further divide active users into sub-categories based on level of engagement: one-time visitors; short-term, long-term, and persistent adopters. Each category reflects the duration of their activity on Mastodon following the migration.
One-time visitors. These academics created an account on Mastodon at the beginning of the Twitter-to-Mastodon migration, posted only within the first week of joining and then abandoned the platform ($n_{days}\leq 7$). This brief engagement suggests a possible exploration of the platform following the migration advocacy, without finding enough value or community to continue their activity.
Short-term adopters. This group includes those who actively used Mastodon for more than one week but did not persist beyond 100 days. They left the platform possibly due to unmet expectations or a stronger pull back to more familiar platforms like Twitter.
Long-term adopters. These users remained active on Mastodon from the time of migration for relatively long time ($100<n_{days}\leq 300$) but have been inactive.
Persistent adopters. These users remained active on Mastodon from initial migration though the end of our analyses ($n_{days}>300$). Their sustained engagement indicates successful integration into the Mastodon community.
3.2.2 Inactive users.
Academics in this category were already using Mastodon before the migration wave from Twitter. Despite their early presence on the platform, they remained inactive during the migration period. One user was inactive during the first month but engaged briefly afterward. Inactive users are not included in subsequent analyses. The distribution of users is given in Table 1. Note that only 35.4% of academic users in our dataset are active and consistent in their use of Mastodon. The majority demonstrate sporadic engagement.
Category Subcategory Count
Active Users One-time visitors 1,183
Short-term adopters 1,624
Long-term adopters 2,065
Persistent adopters 2,665
Inactive Users One-Time Visitors 4
Returned Users 1
Table 1: Number of users, by activity.
We calculate the average following, follower and post counts for each user category, aggregated by month show in Figure 2. Our analysis reveals a sharp increase in all three public metrics at the onset of the migration period, specifically from November 2022 to December 2022. However, beyond this period, the rate of increase for following and follower metrics slows considerably. For one-time, short-term, and long-term users, growth in following and follower counts stagnated beginning in early 2023. As these users were not regularly contributing content, their visibility and potential to attract followers were likewise limited. In contrast, persistent users continued to post regularly. Despite their consistent activity, however, growth in their connections remained slow. This slow growth highlights the challenges faced in establishing and expanding social ties and communities on Mastodon.
Figure 2: Average public metrics and statuses count for active users.
3.3 Field-specific activity for academic accounts on Mastodon
As discussed in Section 2, users self-identified to research disciplines in public lists. We further categorize these disciplines into 9 fields, consistent with existing taxonomies of academic disciplines [1]. Scholars identifying with multiple disciplines and open science researchers were categorized as multi-disciplinary researchers. Inactive users were excluded from subsequent analyses. Disciplinary statistics of the 7,537 academics in our datasets are provided in Figure A5 in the Appendix.
We identify the proportion of one-time, short-term, long-term, and persistent users by field (see Table 2). The highest proportion of persistent users is observed within multidisciplinary researchers. The public metrics aggregated by field and by month are shown in Figure A4 in the Appendix. Researchers in the physical sciences and mathematics have notably higher average number of posts, indicating their consistent and significant contribution to discussions on Mastodon. Scholars in medicine and health sciences exhibit relatively less frequent posting.
Field One Time (%) Short Term (%) Long Term (%) Persistent (%)
Social and Behavioral Sciences 15.7 21.4 27.9 35.0
Arts and Humanities 20.5 19.4 24.8 35.3
Physical Sciences and Mathematics 12.0 21.2 27.8 39.0
Life Sciences 10.7 20.2 41.7 27.4
Education 21.3 26.2 24.2 28.3
Law 17.9 28.3 26.6 27.3
Medicine and Health Sciences 10.8 30.3 29.3 29.6
Multidisciplinary 10.8 18.3 28.8 42.1
Table 2: Distribution of user engagement by academic field.
3.4 Twitter and Mastodon cross-platform analysis
To further investigate the conversations surrounding Mastodon among academics, we leverage data collected from Twitter reflecting conversations about academics on Mastodon (see Section 2). We plot the daily frequency of these tweets over time in Figure 3. Discussions about Mastodon in the academic context experienced a noticeable spike during the first two months of the movement.
Conversations reflect difficulties and concerns during the transition. For example, Mastodon struggled to accommodate the abrupt influx of users. On the peak day of discussion, November 18, 2022:
"Looks like Mastodon isn’t an option anyway. Academic servers aren’t taking new accounts at the moment."
"I have had a Mastodon presence for several months but haven’t used it. I still have a vain hope Twitter will sort itself out, but if not, I am <scholar’s Mastodon account>."
"As you know, I’m over on Mastodon too. Personally I don’t think that I will leave Twitter until and unless Elon finally and truly kills it, but I know that we will connect over on mastodon if he does."
However, this initial surge in discussion was short-lived. The frequency of related tweets sharply declined.
Figure 3: Dynamic frequency of Twitter conversations related to academics migrating to Mastodon.
Users expressed frustration over the platform’s lack of centralized theme and purpose, and its fragmented community structure.
"Vibe-wise, mastodon has *a ton* of different servers already, so there’s more diversity. but the *big* servers that most ppl are on are… kinda academic? kinda techy? very… not fun? it feels like linkedin or facebook groups to a lot of people."
"My humanities folks have largely migrated to mastodon, but my science and public health feeds are still here on Twitter. More’s the pity."
Despite these challenges, ongoing efforts to promote Mastodon indicate a persistent belief in its potential. For instance, in June 2023, one scholar posted,
"Reasons to join us in the fediverse. 500 character limit. free edit button. Lots more #openaccess resources for #teachers, researchers, science communicators, lifelong learners. #STEM #education #openscience #K12 #HigherEd"
Amongst the discussions, we observe concurrent mentions of other platforms such as Bluesky [^3] https://bsky.app/ and Threads [^4] https://www.threads.net/. These have diluted the migration as users explore various platforms for different purposes. For instance,
"I’m curious re categories of people I follow and where they go. IT people seem to have mostly gone to mastodon; librarians ditto; academic leftists stayed here longest, now going to bluesky (prefer commercial sites?); archaeologists have vanished for me completely (threads?)."
"Hey if anyone’s still here on science twitter and wondering where a lot of people have gone, they’re trying out bluesky and liking it. This time it’s real (not like mastodon). #academic"
3.4.1 Case Study: Physical Sciences and Mathematics Scholars.
We observe that scholars in the physical sciences and mathematics were amongst the most engaged on Mastodon. We randomly sampled 100 scholars in these fields who had provided Twitter account information and matched their Mastodon accounts with their Twitter accounts. We recorded whether they are still active on Twitter, and whether they still provide their Mastodon account information in their username, profile description, or pinned tweets. Additionally, we collected follower and following statistics of the sampled scholars on Twitter, all statistics shown in Table 3. We observe that 86% of sampled scholars are still active on Twitter, and among these active users, 62.8% do not provide Mastodon information in easily accessible places, e.g., username, bio, pinned tweets, indicating no signs of migration or intention to migrate. These active Twitter users have accumulated a much higher number of followers, which make it more difficult to give up the platform. The rest of the active users still make their Mastodon information accessible to the community. For instance, one scholar noted:
"Hi ya’ll for those moving over to #Mastodon you can currently find me <Scholar’s Mastodon Account>."
This group of scholars does not personally intend to migrate but aims to maintain connections with those who do intend to fully migrate to Mastodon. In contrast, 5% of the sampled scholars still have accessible but inactive accounts since November 2022, indicating they will not return to Twitter. For example, one scholar’s last post reads,
"I’m leaving Twitter but I’ve been advised not to delete my account in order to avoid possible misuse - so I’ll leave it here as a placeholder. See you at Mastodon!"
Some, 9% of our sampled scholars, chose to lock their accounts as a sign of leaving Twitter, leaving a message in their profile description directing followers to their Mastodon accounts.
Contain Mastodon Info Not Contain Mastodon Info Following Count Follower Count
Active 32 54 Mean: 1593, std: 2684 Mean:11022, std: 30146
Inactive 4 1 Mean: 330, std: 451 Mean: 848, std: 1396
Locked 7 2 Mean: 1551, std:1361 Mean: 1631, std: 1837
Table 3: Engagement statistics of 100 sampled scholars.
In addition to segmentation, many scholars direct their followers to Threads and Bluesky instead of Mastodon, with 2% providing Threads information and 6% providing Bluesky information among the sampled scholars. This trend indicates that Mastodon faces significant competition from other emerging platforms, making it even more challenging for Mastodon to become a preferred alternative for academic discourse.
In summary, we observe that the majority of scholars are either solely using Twitter or using both Twitter and Mastodon simultaneously. Comparing the follower and following counts across these platforms, Mastodon hosts a smaller community. Competition from other emerging social media platforms further limits Mastodon’s growth, indicating that scholars are exploring various alternatives rather than committing to Mastodon.
We cross-mapped Twitter activity levels with Mastodon role categories, as shown in Table 4. Our analysis reveals that 55% of the sampled scholars are relatively active on both Mastodon (long-term or persistent users) and Twitter (active users). Among those who are inactive on Twitter or have locked their accounts, 71% are long-term or persistent users of Mastodon. This indicates that while Mastodon has some loyal users, their numbers are relatively small compared to the entire academic community.
One_Time Short_Term Long_Term Persistent
Active 12 19 28 27
Inactive 1 1 2 1
protected 0 2 1 6
Table 4: Mapping of Twitter (Active, Inactive, Protected) and Mastodon (One Time, Short Term, Long Term, Persistent) activity levels.
4 Discussion and Conclusion
Our analysis of professional communities migrating between social media platforms highlights the unique challenges these users face during such transitions. While decentralized platform offered a promising alternative, the established communities and clear audience engagement on Twitter proved too significant to overcome. The fragmented communities on Mastodon lacked the cohesiveness found on Twitter. Our results show that the initial surge in Mastodon sign-ups did not translate into sustained user engagement.
Among the users we tracked, the majority failed to maintain the same level of activity. This could be attributed to steep learning curve compared to twitter as well as competition from more user-friendly platforms like Bluesky and Threads. These alternatives might have drawn away potential users exacerbating Mastodonś difficulties in becoming a central hub for professionals communities. Additionally, those who continued to post consistently on Mastodon, did not receive same level of engagement as on Twitter which is crucial for sustaining professional discourse.
Our study highlights that without a good alternative, professionals will naturally gravitate back to familiar platforms where their communities are already established through years. Future efforts that support migrations to decentralized platforms must address these core issues, foster community building by reducing friction, and considering network effects to retain user engagement. While the idea of decentralization of social media platforms holds a significant promise, practical challenges persist that require more than just technical solutions. Successful transitioning involves creating communities that rival established social media networks and user experiences of platforms such as Twitter.
{credits}
[…]
There is no competing interest.

### v2

@@
 The Failed Migration of Academic Twitter
-Abstract
-Following change in Twitter’s ownership and subsequent changes to content moderation policies, many in academia looked to move their discourse elsewhere and migration to Mastodon was pursued by some.
+Abstract.
+Following changes in Twitter’s ownership and subsequent changes to content moderation policies, many in academia looked to move their discourse elsewhere and migration to Mastodon was pursued by some.
 Our study looks at the dynamics of this migration.
 Utilizing publicly available user account data, we track the posting activity of academics on Mastodon over a one year period.
+We also gathered follower-followee relationships to map internal networks, finding that the subset of academics who engaged in migration were well-connected.
+However, this strong internal connectivity was insufficient to prevent users from returning to Twitter/X.
 Our analyses reveal significant challenges sustaining user engagement on Mastodon due to its decentralized structure as well as competition from other platforms such as Bluesky and Threads.
-The movement lost momentum after an initial surge of enthusiasm as most users did not maintain their activity levels, and those who did faced lower levels of engagement compared to Twitter.
-Our findings highlight the challenges involved in transitioning professional communities to decentralized platforms, emphasizing the need for focusing on migrating social connections for long-term user engagement.
-Keywords:
-Social Media Mastodon Twitter Social Movement Computational Social Science Decentralized Platforms.
-**footnotetext: These authors contributed equally to this work.
-1 Introduction
+The movement lost momentum after an initial surge of enthusiasm where the main network was fully established as most users did not maintain their activity levels, and those who did faced lower levels of engagement.
+Our findings highlight the challenges involved in transitioning professional communities to decentralized platforms, emphasizing the need for focus on community building for long-term user engagement.
+Social Media, Mastodon, Twitter, Social Movement, Computational Social Science, Decentralized Platforms
+copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2024; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Social networking sitesccs: Human-centered computing Social mediaccs: Human-centered computing Computer supported cooperative workccs: Human-centered computing Social networks
+1. Introduction
 Social media platforms have become essential outlets for communication, information dissemination, and professional networking.
-Among these platforms, Twitter (now known as “X”) has gained significant popularity, often referred to as virtual town square for its role facilitating public discourse [10].
-Academics and journalists populate sub-communities on Twitter and use the platform for professional communication [7, 9].
-Academics have extensively used Twitter to share research findings, announce academic events such as workshops and conferences, and contribute to wider scholarly discourse [11, 8].
-However, after Elon Musk’s acquisition of the company and controversial changes to content moderation policies, many on so-called Academic Twitter sought alternatives [2].
-Among academics, Mastodon, a decentralized social media platform, emerged as a popular alternative with some creating servers themed around their topics of interest.
-Unlike traditional platforms, Mastodon operates on a federated model where independent servers interact through a common protocol, allowing users to join different instances based on their preferences while still communicating across the entire network [14].
-This structure aims to reduce centralized control over internet services by distributing data and services across multiple nodes, enhancing privacy, security, and resilience against censorship and corporate control [12].
-Yet despite these advantages, a large-scale measurement study of Mastodon revealed that this decentralized structure also leads to lower user engagement compared to centralized platforms like Twitter, where network effects and a larger, more active user base drive higher engagement levels [12].
+Among these platforms, Twitter (now known as “X”) has gained significant popularity and has been likened to a virtual town square for its role facilitating public discourse (Mascaro and Goggins, 2012).
+Academics and journalists populate sub-communities on Twitter and use the platform for professional communication (Jurkowitz and Gottfried, 2022; Malik et al., 2019).
+Academics have extensively used Twitter to share research findings, announce academic events such as conferences, and contribute to wider scholarly discourse (Mohammadi et al., 2018; Ebner, 2013; Mahrt et al., 2014).
+However, after Elon Musk’s acquisition of the company and controversial changes to content moderation policies, many on so-called Academic Twitter sought alternatives (Cava et al., 2023).
+Among academics, Mastodon, a decentralized social media platform, emerged as a popular alternative with some creating servers themed around their areas of interest.
+Unlike traditional platforms, Mastodon operates on a federated model where independent servers interact through a common protocol, allowing users to join different instances based on their preferences while still communicating across the entire network (Zignani et al., 2018).
+This structure aims to reduce centralized control over internet services by distributing data and services across multiple nodes, enhancing privacy, security, and resilience against censorship and corporate control (Raman et al., 2019).
+Yet despite these advantages, a large-scale measurement study of Mastodon revealed that this decentralized structure also leads to lower user engagement compared to centralized platforms like Twitter, where network effects and a larger, more active user base drive higher engagement levels (Raman et al., 2019).
 The migration of users between social media platforms has been a topic of interest among researchers, particularly in the context of changes in platform policies, user experiences, and factors that influence user decisions and the implications for online communities.
-Social media movements often occur in response to significant changes in platform policies, ownership, user dissatisfaction or perceived deficiencies in a platform[13].
-Qualitative analyses, including interviews and surveys of users migrating from platforms like Fandom to alternatives such as Tumblr, reveal that policy and value-based aspects are primary catalysts in the migration process [3].
-Additionally, research has shown that user migration is influenced by factors such as perceived platform attractiveness and peer influence [5].
-Study of users migrating from Twitter to Mastodon post-Elon Musk’s acquisition indicated that social influence, network structure, commitment, and shared identity significantly drive collective behavior changes in online communities [2].
-Analysis showed Twitter-to-Mastodon migration was characterized by low Twitter account deletion rates, concentration of users in large Mastodon instances, varied posting behaviors across platforms [4].
-A similar study on Twitter-to-Mastodon migration revealed lower user engagement and posting activity on Mastodon, highlighting challenges in retaining active users due to its decentralized architecture [6].
-Despite prior work, there remains a gap in the research concerning specific professional communities, as previous work has focused on the general population of social media platform users.
+Social media movements often occur in response to significant changes in platform policies, ownership, user dissatisfaction or perceived deficiencies in a platform (Xu et al., 2014).
+Qualitative analyses, including interviews and surveys of users migrating from platforms like Fandom to alternatives such as Tumblr, reveal that policy and value-based aspects are primary catalysts of change (Fiesler and Dym, 2020).
+Additionally, research has shown that user migration is influenced by perceived platform attractiveness and peer effects (Hou and Shiau, 2020).
+A study of users migrating from Twitter to Mastodon post-Musk’s acquisition indicated that social influence, network structure, commitment, and shared identity significantly drive collective behavior changes in online communities (Cava et al., 2023).
+Analyses showed Twitter-to-Mastodon migration was characterized by low Twitter account deletion rates, concentration of users in large Mastodon instances, varied posting behaviors across platforms (He et al., 2023).
+A similar study on Twitter-to-Mastodon migration revealed lower user engagement and posting activity on Mastodon, highlighting challenges in retaining active users due to its decentralized architecture (Jeong et al., 2023).
+Despite prior work characterizing the general population of platform users, there remains a gap in targeted research concerning specific professional communities.
 In this study, we address this gap by examining the migration patterns and behaviors of academic users, providing a more nuanced understanding of how these particular communities navigate platform transitions.
+Given the significant role that Twitter plays in an academic’s professional life (Knight and Kaye, 2016; Klar et al., 2020; Stewart, 2015), academics represent a uniquely valuable group for studying the migratory dynamics.
 Our study addresses the following research questions:
 RQ1: To what extent were academics engaged in the migration from Twitter to Mastodon, and how can the success or failure of this migration be assessed?
-RQ2: Which academic disciplines have demonstrated the highest levels of persistence on Mastodon, and what factors characterize their migration experiences?
-Our findings indicate that majority of the academic researchers moved to Mastodon failed to maintain their posting activity and follower engagement.
-These results underscore the importance of not only transferring accounts but also migrating social connections to ensure the success of social media transitions.
-2 Data
-We utilized a public Github repository containing lists of academics on Mastodon [^1] https://github.com/nathanlesage/academics-on-mastodon?tab=readme-ov-file.
-These lists were curated by researchers from 50 distinct disciplines, who volunteered to be added to the lists.
-Each list corresponds to a specific research discipline, containing account details of researchers within that research area.
-If a list for a particular discipline does not exist, one can start a new list and provide a corresponding form.
-This form allows researchers to add themselves to the relevant list based on the discipline they identify with.
-While the database continues to expand, for the purpose of tracking, we focused only on scholars whose Mastodon accounts were already in the repository at the beginning of November 2022, noting the Twitter-to-Mastodon migration began in late October 2022.
-In total, we tracked the Mastodon profiles and public metrics of 7542 academic users using a customized Mastodon API.
-Data was collected weekly from November 2022 to October 2023 allowing for temporal analysis of the user activity and engagement over the span of a year.
-For cross-platform analysis, we also collected tweets via Zeeschuimer [^2] https://github.com/digitalmethodsinitiative/zeeschuimer using the query "mastodon (researcher OR researchers OR academic OR academics OR professor OR professors OR scholar OR scholars) lang:en until:2023-11-01 since:2022-10-01".
-Since many tweets were included because the username contained the word "mastodon" (as some users use their Mastodon account name as their Twitter username, we further filtered out these tweets).
+RQ2: Which academic fields have demonstrated the highest levels of persistence on Mastodon, and what factors characterize their migration experiences?
+Our findings indicate that majority of academic researchers who moved to Mastodon failed to maintain their posting activity and follower engagement over long term.
+Although migrated academic accounts formed strongly-connected internal networks, these connections were insufficient to sustain user engagement over time.
+Of note, however, communities that established field specific server instances had higher user retention rate compared to those on general-purpose servers.
+Cross platform analysis of most recent Twitter activity showed that most users returned to Twitter.
+These results underscore the challenges involved and the potential opportunities to build and foster sustainable communities at scale on new platforms.
+2. Related Work
+2.1. Characteristics of Mastodon
+The decentralized structure Mastodon and similar platforms leads to unique user engagement dynamics that are different from centralized platforms.
+Research suggests that users are drawn to Mastodon for the ease of protecting their information from data mining and the absence of algorithms shaping their feeds (Lee and Wang, 2023).
+However, this can also make content discovery more difficult (Zignani et al., 2019).
+The decentralized structure encourages strong, tight-knit communities within individual server instances but fewer connections across instances (Zignani et al., 2018).
+These characteristics pose challenges for migration as well.
+2.2. Cross-platform migration
+Social media movements often occur in response to significant changes in platform policies, ownership, user dissatisfaction or perceived deficiencies in a platform (Xu et al., 2014).
+The migration of users between social media platforms has been a topic of interest among researchers, particularly in the context of changes in platform policies, user experiences, and factors that influence user decisions and the implications for online communities’ retention.
+Qualitative analyses, including interviews and surveys of users migrating from platforms like Fandom to alternatives such as Tumblr, reveal that policy and value-based aspects are primary catalysts in the migration process (Fiesler and Dym, 2020).
+Additionally, research has shown that user migration is influenced by factors such as perceived platform attractiveness and peer influence (Hou and Shiau, 2020).
+Study of users migrating from Twitter to Mastodon post-Elon Musk’s acquisition indicated that social influence, network structure, commitment, and shared identity significantly drive collective behavior changes in online communities (Cava et al., 2023).
+Analysis showed Twitter to Mastodon migration is characterized by low Twitter account deletion rates, concentration of users in large Mastodon instances, varied posting behaviors across platforms (He et al., 2023).
+A similar study on Twitter-to-Mastodon migration reveals lower user engagement and posting activity on Mastodon, highlighting challenges in retaining active users due to its decentralized architecture (Jeong et al., 2023).
+Our work looks more deeply at the specific case of the relatively well-synchronized and coordinated movement of communities of academic users across platforms.
+3. Data
+Note on terminology: Twitter rebranded to X in July 2023, which occurred midway through our data collection. For simplicity, we use the terminology Twitter and tweets throughout the paper.
+Our analyses are primarily based on a public Github repository containing lists of academics on Mastodon [^1] https://github.com/nathanlesage/academics-on-mastodon.
+These lists were curated by researchers from 50 distinct disciplines, who voluntarily added their own names and accounts.
+Each list corresponds to a research discipline, containing account details of researchers within that area, their Mastodon IDs, and a short description of their research.
+Some lists also included each user’s Twitter handle.
+While the database continues to expand, for the purpose of longitudinal analysis, we focused only on scholars whose Mastodon accounts were already in the repository at the beginning of November 2022, noting the Twitter-to-Mastodon migration began in late October 2022.
+That is, the users in our dataset were part of a coordinated, synchronized migration effort in November 2022.
+In total, we collected the Mastodon profiles and corresponding public metrics, e.g., number of followers, following, timestamps of most recent status of 7,542 academic user accounts using a customized Mastodon API [^2] https://github.com/halcy/Mastodon.py.git.
+Data was collected weekly from November 2022 to October 2023 allowing for temporal analysis of user activity and engagement over the span of a year.
+In addition, we collected the most recent Mastodon data related to follower and following accounts for each scholar to conduct network analyses based on fields of study and associated servers.
+For cross-platform analyses, we used self-reported Twitter handles to gather scholars’ profile information.
+Of 7,542 Mastodon scholars, a subset of 3,131 scholars accurately provided Twitter handles suitable for data collection.
+Using the Twitter (“X”) Basic API, we retrieved profile information, most recent tweet IDs, and engagement metrics.
+Additionally, we used the Twitter ID collection endpoint to extract the posting time of their most recent tweet.
+Finally, to understand discussions surrounding Mastodon in academic context, we collected tweets during the one year window via Zeeschuimer [^3] https://github.com/digitalmethodsinitiative/zeeschuimer using the boolean search “mastodon (researcher OR researchers OR academic OR academics OR professor OR professors OR scholar OR scholars) lang:en until:2023-11-01 since:2022-10-01”.
+Since many tweets were included because the username contained the word “mastodon” (as some users use their Mastodon account name as their Twitter username) we further filtered out these tweets.
 We retained only those that contained relevant keywords in the main body of the tweet, resulting in a total of 1,497 tweets.
-3 Results
-3.1 Attrition of academic accounts on Mastodon
-Figure 1 displays a count heat map illustrating the relationship between “last status time" and “record time”, aggregated by month.
-The “last status time” refers to the time of the user’s most recent post, while “record time" indicates the time at which the data was collected.
+Figure 1. Waterfall plot showing number of monthly active users (active=at least one post per month).
+4. Results
+Following, we present findings from multiple analyses conducted to understand the dynamics of the migration.
+First, we used the longitudinal data from Mastodon to understand user attrition and engagement trends over the year.
+Then, we used network analysis to understand the connectivity and distribution of academics across Mastodon servers.
+Next, we did a cross-platform analysis using the subset of users who provided Twitter handles comparing their activity patterns on Twitter and Mastodon.
+Finally, we ran exploratory analyses on Twitter posts related to the migration to understand user attitudes towards migration.
+4.1. Longitudinal Analysis
+4.1.1. Attrition of academic accounts on Mastodon
+Figure 1 gives a count waterfall plot illustrating the change in number of active users aggregated by month and a line plot showing the monthly retention rate based on the previous month.
+In this context, we considered a user to be active if they made at least a post in a given month.
 In the initial month of data collection (November 2022), there were 7,505 active scholars who had posted within that same month.
-This number experienced a gradual decline over the subsequent seven months, showing a minor resurgence in July 2023 before continuing its descent, ultimately reaching 2,398 by October 2023.
-The figure also shows that 1,237 scholars were active in November 2022 but then not again throughout the remainder of our data collection.
-We observe a significant rate of attrition; each month, approximately 10% to 20% of scholars in our set stopped actively using the platform.
-Figure 1: Heat map of user activity. Last status vs. record time (aggregated by month).
-3.2 Engagement with academic accounts on Mastodon
-Given the central role of engagement in user activity, we look at the engagement with migrated academic accounts on Mastodon.
-To do, so we first characterize accounts based on their activity on the platform.
-3.2.1 Active users.
+[^4] There were 1, 4, and 32 scholars had their last activity in April, May, and October 2022, respectively, which fell outside our collection window.
+These scholars were excluded from the waterfall plot but included in subsequent analysis.
+Over the next seven months, the number of active scholars experienced a decline, showing a minor resurgence in July 2023 before continuing its descent, ultimately reaching 2,398 by October 2023.
+We observe a subtantial rate of attrition; each month, approximately 10% to 20% of scholars in our set stopped actively using the platform.
+4.1.2. Engagement with academic accounts on Mastodon
+Given the central role of engagement in user activity, we look at the engagement with migrated academic accounts on Mastodon. To do so, we first characterize accounts based on their activity.
+Active users.
 We define active users as scholars who engaged on Mastodon during the initial Twitter-to-Mastodon migration, which began in late October 2022.
@@
 Their sustained engagement indicates successful integration into the Mastodon community.
-3.2.2 Inactive users.
+4.1.3. Inactive users.
 Academics in this category were already using Mastodon before the migration wave from Twitter.
@@
 The majority demonstrate sporadic engagement.
-Category Subcategory Count
-Active Users One-time visitors 1,183
+User Category Subcategory Count
+Active One-time visitors 1,183
 Short-term adopters 1,624
@@
 Persistent adopters 2,665
-Inactive Users One-Time Visitors 4
+Inactive One-Time Visitors 4
 Returned Users 1
-Table 1: Number of users, by activity.
+Table 1. Number of users, by activity.
 We calculate the average following, follower and post counts for each user category, aggregated by month show in Figure 2.
@@
 This slow growth highlights the challenges faced in establishing and expanding social ties and communities on Mastodon.
-Figure 2: Average public metrics and statuses count for active users.
-3.3 Field-specific activity for academic accounts on Mastodon
-As discussed in Section 2, users self-identified to research disciplines in public lists.
-We further categorize these disciplines into 9 fields, consistent with existing taxonomies of academic disciplines [1].
+The persistent users’ continued efforts to engage and post suggest a commitment to the platform, yet the limited increase in followers indicates potential barriers in user retention and community building.
+Figure 2. Average public metrics and statuses count for active users.
+4.1.4. Field-specific activity differences among academic accounts on Mastodon
+As discussed in Section 3, users self-identified their research disciplines in public lists.
+We aggregate these disciplines into 9 broad fields, consistent with existing taxonomies of academic disciplines (bepress., 2019).
 Scholars identifying with multiple disciplines and open science researchers were categorized as multi-disciplinary researchers.
 Inactive users were excluded from subsequent analyses.
-Disciplinary statistics of the 7,537 academics in our datasets are provided in Figure A5 in the Appendix.
+Disciplinary statistics of the 7,537 academics in our datasets are provided in Table A1 in the Appendix.
 We identify the proportion of one-time, short-term, long-term, and persistent users by field (see Table 2).
@@
 Scholars in medicine and health sciences exhibit relatively less frequent posting.
-Field One Time (%) Short Term (%) Long Term (%) Persistent (%)
-Social and Behavioral Sciences 15.7 21.4 27.9 35.0
-Arts and Humanities 20.5 19.4 24.8 35.3
-Physical Sciences and Mathematics 12.0 21.2 27.8 39.0
-Life Sciences 10.7 20.2 41.7 27.4
-Education 21.3 26.2 24.2 28.3
-Law 17.9 28.3 26.6 27.3
-Medicine and Health Sciences 10.8 30.3 29.3 29.6
-Multidisciplinary 10.8 18.3 28.8 42.1
-Table 2: Distribution of user engagement by academic field.
-3.4 Twitter and Mastodon cross-platform analysis
+Field One Time (%) Short Term (%) Long Term (%) Persistent (%) Total Percent In-Degree Out-Degree
+Social & Behavioral Sciences 15.7 21.4 27.9 35.0 2918 38.72 98881 91958
+Arts & Humanities 20.5 19.4 24.8 35.3 1522 20.19 38692 39986
+Physical Sciences & Mathematics 12.0 21.2 27.8 39.0 851 11.29 11453 11980
+Life Sciences 10.7 20.2 41.7 27.4 84 1.11 338 648
+Education 21.3 26.2 24.2 28.3 244 3.24 2249 4171
+Law 17.9 28.3 26.6 27.3 587 7.79 21664 27450
+Medicine & Health Sciences 10.8 30.3 29.3 29.6 287 3.81 4292 4923
+Multidisciplinary 10.8 18.3 28.8 42.1 1040 13.80 56371 52818
+Business 25 25 25 25 4 0.05 60 66
+Table 2. Distribution of user engagement and node degree statistics by academic field.
+4.2. Mastodon Network Analyses
+To understand the connectivity within and between different fields, we collected the complete follower-followee data of all academic accounts on Mastodon, and constructed an internal connection network within each academic group, as well as a network aggregated by field of study.
+4.2.1. Internal Network Aggregated by Field
+Figure 3 shows the network connections across different fields.
+In this network, node size represents the number of academics in each field, while edge width indicates the frequency of connections between the user accounts from different fields.
+Given the nature of the field, certain fields such as Social and Behavioral Sciences are strongly connected to Multi-disciplinary field.
+Our analysis reveals that the academic migration has led to the formation of a well-connected internal network, suggesting that researchers actively attempted to establish a new community on Mastodon.
+However, despite these coordinated efforts, the migration still faced significant challenges, as reflected through the low retention rates shown in Section 4.1.
+4.2.2. Internal Networks aggregated by Server
+Prior studies have shown that the Mastodon’s server instance based structure makes it easier for users to discover and connect with others within the same instance than outside instance.
+This leads to stronger within instance ties compared to weaker links across instances (Zignani et al., 2019).
+Therefore, we analyzed the follower-followee network aggregated by Mastodon instances, shown in Figure 4.
+The academic users were spread across a total of 342 servers, with 9 server instances being isolates without any connecting edges to other instances, which were removed for clarity.
+Table A2 in the Appendix shows network centrality metrics for each server and the distribution of the users across different servers.
+It also includes the official description of the server instances, indicating whether they are customized to a specific group of scholars.
+During the migration, different disciplines used different strategies: some used general-purpose servers like mastodon.social which host a wide variety of content beyond just academic content, while others took a coordinated approach on setting up their field specific server instances such as infosec.exchange and journa.host.
+Figure 4 shows the follower-followee connections across different servers showing certain general purpose instances such as mastodon.social are popular among the tracked academic users (containing 21.3% of the total) than other general purpose servers like mas.to.
+Figure 3. internal field network (size: number of scholars in each field; edge weights: frequency of connections; edge color: same as the source).
+4.2.3. Case study: Mastodon Community of Information Security
+We also examined a discipline-specific server, unique to decentralized social platforms, and its user retention rate using a case study of the Information Security discipline.
+We chose this discipline based on our field-level engagement analysis shown in Figure A1 in Appendix, in the Appendix, which indicates that scholars in physical sciences and mathematics were among the most active on Mastodon.
+To further understand the impact of server instance choice on community formation, we visualized their internal network, highlighting the top five servers where these scholars are based.
+Figure A2 shows the internal follower-followee network for physical science and mathematics field, highlighting the top five servers from a total of 33 servers that these scholars use.
+Interestingly, scholars from Information Security formed a strongly connected community on a custom discipline specific server, unlike other disciplines that used general servers such as mastodon.social or fediscience.org.
+To explore how a discipline specific server could influence user retention and engagement, we isolated the internal network of accounts within Information Security discipline.
+Figure 5 shows the internal network of the discipline, where most scholars congregated on a single server: infosec.exchange (highlighted in red).
+We compare the retention roles of scholars from the discipline with those from a discipline where users were concentrated in a general-purpose server and other sub-fields of similar size.
+We see a significantly high percentage of persistent users for information security scholars (56.8%) in comparison to other comparable disciplines.
+This suggests that establishing field- or discipline-specific communities or servers could foster a stronger sense of belonging, which may help retain scholars who have migrated, reducing the likelihood of their return to Twitter or seeking alternative platforms.
+Figure 4. internal server network (removed isolated servers, red are servers with more than 100 scholars).
+Figure 5. internal follower-followee network for Information Security community (removed isolates).
+Category Info. Sec. Education Communication & Media Studies
+Count Percent Count Percent Count Percent
+One-Time 23 8.7 52 21.3 41 15.5
+Short-Term 31 11.7 64 26.2 53 20.0
+Long-Term 60 22.7 59 24.2 81 30.6
+Persistent 150 56.8 69 28.3 90 34.0
+Total 264 100.0 244 100.0 265 100.0
+Table 3. User retention breakdown for Information Security scholars compared to the sub-fields that are similar in size.
+4.3. Twitter-Mastodon Cross-platform Analysis
+Since our attrition analysis showed that only a few users remained active on Mastodon, we conducted a cross-platform analysis to analyze their account state on Twitter.
+Approximately 50% of the scholars voluntarily provided their Twitter usernames, which were accurate and traceable to their Twitter profile pages.
+Using the Official Twitter Basic API, we collected the profile information of these accounts.
+Despite some account deletions, we successfully retrieved 3,131 Twitter accounts that could be matched to Mastodon accounts.
+Next, we retrieved the most recent posting times of these Twitter accounts to assess their activity status.
+Based on their posting behavior, we categorized the Twitter account activity into the following statuses:
+Locked: Scholar’s account is locked down, meaning their posts are private and cannot be accessed.
+Inactive: Scholar’s latest tweet was posted before December 31st, 2022, and their posts were made before this date, indicating no recent activity since the start of the year.
+Semi-active: Scholar’s latest tweet was posted between December 31st, 2022 and November 1, 2023.
+This indicates some activity during the year but no recent posts after November, 2023, the last timestamp of our Mastodon tracking.
+Active: Scholar’s latest tweet was posted after November 1st, 2023, indicating continuous activity and engagement on the platform.
+Using the above categorization, we calculated the count and connection statistics for each category, and all statistics are presented in Table 4 (left).
+We observe that 79.72% of scholars are still active on Twitter, indicating no signs of migration or intention to fully migrate.
+These active Twitter users have accumulated a much higher number of followers, which make it more difficult to give up the platform.
+Some active users still make their Mastodon information accessible to the community.
+For instance, one scholar noted:
+”Hi ya’ll for those moving over to #Mastodon you can currently find me ¡Scholar’s Mastodon Account¿.”
+This group of scholars does not personally intend to migrate but aims to maintain connections with those who do intend to fully migrate to Mastodon.
+Furthermore, 7.03% of the scholars are semi-active users on Twitter.
+They did not stop using Twitter during the migration, but have reduced their presence gradually during a one-year grace period, potentially transitioning away from Twitter entirely.
+In contrast, 2.91% of the scholars still have accessible but inactive accounts since November 2022, indicating they will not return to Twitter. For example, one scholar’s last post reads,
+”I’m leaving Twitter but I’ve been advised not to delete my account in order to avoid possible misuse - so I’ll leave it here as a placeholder. See you at Mastodon!”
+Approximately 10.35% of the scholars chose to lock their Twitter accounts, often signaling their departure from the platform by leaving a message in their profile directing followers to their Mastodon accounts.
+Figure 6. Connection metrics mapping between mastodon and Twitter.
+Count Following Count Follower Count One-time Short-term Long-term Persistent
+Active 2496 Mean:1989.07, std:3087.68 Mean:13208.76 , std:108495.54 374 578 684 860
+Semi-active 220 Mean:1137.56 , std:1263.87 Mean:4467.28, std:19152.90 39 42 43 96
+Inactive 91 Mean:814.56 , std:1062.39 Mean:2401.46 , std:5203.50 12 19 21 39
+Locked 324 Mean:1859.72 , std:2300.91 Mean:6347.10 , std:31583.98 51 51 82 140
+Table 4. Engagement statistics of 3,131 scholars and mapping of Twitter (Active, Inactive, Protected) and Mastodon (One Time, Short Term, Long Term, Persistent) activity levels.
+We cross-mapped Twitter activity levels with Mastodon role categories, as shown in Table 4 (right).
+Our analysis reveals that 49.31% of the sampled scholars are relatively active on both Mastodon (long-term or persistent users) and Twitter (active users).
+Among those who are inactive on Twitter or have locked their accounts, 67.95% are long-term or persistent users of Mastodon.
+However, this group of dedicated Mastodon users represents only 9.02% of the total scholars included in the cross-platform analysis.
+This suggests that while there is a core group of loyal users who have fully migrated to Mastodon, they make up a relatively small portion of the overall academic community.
+We also plotted the pairwise follower and following counts for the Mastodon and Twitter accounts of the same 3,131 scholars, as shown in Figure 6.
+The difference in connection scale is clear: while scholars who migrated to Mastodon made efforts to establish a new community, it is nowhere near the scale of connections they built over the years on Twitter.
+Additionally, we plotted the Twitter account creation date against the total number of tweets, displayed in Figure A3.
+The color of the points represents the total number of likes received by each account.
+A great proportion of scholars created their Twitter accounts between 2008 and 2012, indicating that the deep, long-standing networks, high levels of engagement, and rich posting histories developed on Twitter are not easily replicable or discarded.
+This entrenched presence highlights the difficulty of transitioning to a new platform like Mastodon without sacrificing years of accumulated influence and community ties.
+4.3.1. Alternatives to Mastodon
+In addition to Mastodon, other centralized alternatives to Twitter have emerged in recent years, and many academics explored these venues.
+Unlike Mastodon, these platforms provided familiar interfaces and ease of transitioning.
+Those present on these alternative platforms often added their account information in their Twitter account descriptions to make it easier for others to discover and connect with them on these new platforms.
+To understand the extent to which academics were exploring the alternatives,we analyzed their usernames and user descriptions for mentions of these platforms.
+Specifically, we focused on mentions of Bluesky [^5] https://bsky.app/ and Threads [^6] https://www.threads.net/ using keyword matching as well as user name pattern matching.
+For Mastodon and Bluesky, we also included scholars who use decentralized platform-style usernames as their Twitter usernames (e.g., ¡username¿@¡servername¿).
+Figure 7 shows an UpSet plot illustrating the number of users who explicitly mention each platform or a combination of platforms.
+In addition to 626 scholars mentioning Mastodon, 308 and 69 users also reference BlueSky and Threads, respectively, with some scholars providing account information for multiple platforms.
+This indicates that while these scholars have accounts on Mastodon, only about 20% actively direct their Twitter connections to it.
+The trend highlights that Mastodon is not the sole option for scholars migrating away from Twitter and faces significant competition from other emerging platforms.
+This competition further complicates Mastodon’s ability to establish itself as the primary platform for academic discourse.
+Figure 7. Mentions of Platforms in Username and Profile Description.
+In summary, we observe that the majority of scholars are either solely using Twitter or using both Twitter and Mastodon simultaneously.
+Comparing the follower and following counts across these platforms, Mastodon hosts a smaller community.
+Competition from other emerging social media platforms further limits Mastodon’s growth, indicating that scholars are exploring various alternatives rather than committing to Mastodon.
+4.4. Analysis of Tweets on Mastodon Migration
 To further investigate the conversations surrounding Mastodon among academics, we leverage data collected from Twitter reflecting conversations about academics on Mastodon (see Section 3).
-We plot the daily frequency of these tweets over time in Figure 3.
+We plot the daily frequency of these tweets and engagement counts over time aggregated by week in Figure 8.
 Discussions about Mastodon in the academic context experienced a noticeable spike during the first two months of the movement.
+After the migration period, although conversations were less frequent, they received more attention (likes), indicating that the audience and connections of academics concerned about Mastodon have largely remained active on Twitter.
 Conversations reflect difficulties and concerns during the transition. For example, Mastodon struggled to accommodate the abrupt influx of users. On the peak day of discussion, November 18, 2022:
-"Looks like Mastodon isn’t an option anyway. Academic servers aren’t taking new accounts at the moment."
-"I have had a Mastodon presence for several months but haven’t used it. I still have a vain hope Twitter will sort itself out, but if not, I am <scholar’s Mastodon account>."
-"As you know, I’m over on Mastodon too.
-Personally I don’t think that I will leave Twitter until and unless Elon finally and truly kills it, but I know that we will connect over on mastodon if he does."
+”I have had a Mastodon presence for several months but haven’t used it. I still have a vain hope Twitter will sort itself out, but if not, I am ¡scholar’s Mastodon account¿.”
+”As you know, I’m over on Mastodon too.
+Personally I don’t think that I will leave Twitter until and unless Elon finally and truly kills it, but I know that we will connect over on mastodon if he does.”
 However, this initial surge in discussion was short-lived. The frequency of related tweets sharply declined.
-Figure 3: Dynamic frequency of Twitter conversations related to academics migrating to Mastodon.
+Figure 8. Dynamic frequency of Twitter conversations related to academics migrating to Mastodon and corresponding engagement metrics (connection and like counts), aggregated by week.
 Users expressed frustration over the platform’s lack of centralized theme and purpose, and its fragmented community structure.
-"Vibe-wise, mastodon has *a ton* of different servers already, so there’s more diversity.
+”Vibe-wise, mastodon has *a ton* of different servers already, so there’s more diversity.
 but the *big* servers that most ppl are on are… kinda academic?
@@
 very… not fun?
-it feels like linkedin or facebook groups to a lot of people."
-"My humanities folks have largely migrated to mastodon, but my science and public health feeds are still here on Twitter. More’s the pity."
+it feels like linkedin or facebook groups to a lot of people.”
+”My humanities folks have largely migrated to mastodon, but my science and public health feeds are still here on Twitter. More’s the pity.”
 Despite these challenges, ongoing efforts to promote Mastodon indicate a persistent belief in its potential. For instance, in June 2023, one scholar posted,
-"Reasons to join us in the fediverse.
+”Reasons to join us in the fediverse.
 500 character limit.
@@
 Lots more #openaccess resources for #teachers, researchers, science communicators, lifelong learners.
-#STEM #education #openscience #K12 #HigherEd"
-Amongst the discussions, we observe concurrent mentions of other platforms such as Bluesky [^3] https://bsky.app/ and Threads [^4] https://www.threads.net/.
+#STEM #education #openscience #K12 #HigherEd”
+Amongst the discussions, we again observe concurrent mentions of other platforms such as Bluesky and Threads.
 These have diluted the migration as users explore various platforms for different purposes.
 For instance,
-"I’m curious re categories of people I follow and where they go.
-IT people seem to have mostly gone to mastodon; librarians ditto; academic leftists stayed here longest, now going to bluesky (prefer commercial sites?); archaeologists have vanished for me completely (threads?)."
-"Hey if anyone’s still here on science twitter and wondering where a lot of people have gone, they’re trying out bluesky and liking it. This time it’s real (not like mastodon). #academic"
-3.4.1 Case Study: Physical Sciences and Mathematics Scholars.
-We observe that scholars in the physical sciences and mathematics were amongst the most engaged on Mastodon.
-We randomly sampled 100 scholars in these fields who had provided Twitter account information and matched their Mastodon accounts with their Twitter accounts.
-We recorded whether they are still active on Twitter, and whether they still provide their Mastodon account information in their username, profile description, or pinned tweets.
-Additionally, we collected follower and following statistics of the sampled scholars on Twitter, all statistics shown in Table 3.
-We observe that 86% of sampled scholars are still active on Twitter, and among these active users, 62.8% do not provide Mastodon information in easily accessible places, e.g., username, bio, pinned tweets, indicating no signs of migration or intention to migrate.
-These active Twitter users have accumulated a much higher number of followers, which make it more difficult to give up the platform.
-The rest of the active users still make their Mastodon information accessible to the community.
-For instance, one scholar noted:
-"Hi ya’ll for those moving over to #Mastodon you can currently find me <Scholar’s Mastodon Account>."
-This group of scholars does not personally intend to migrate but aims to maintain connections with those who do intend to fully migrate to Mastodon.
-In contrast, 5% of the sampled scholars still have accessible but inactive accounts since November 2022, indicating they will not return to Twitter.
-For example, one scholar’s last post reads,
-"I’m leaving Twitter but I’ve been advised not to delete my account in order to avoid possible misuse - so I’ll leave it here as a placeholder. See you at Mastodon!"
-Some, 9% of our sampled scholars, chose to lock their accounts as a sign of leaving Twitter, leaving a message in their profile description directing followers to their Mastodon accounts.
-Contain Mastodon Info Not Contain Mastodon Info Following Count Follower Count
-Active 32 54 Mean: 1593, std: 2684 Mean:11022, std: 30146
-Inactive 4 1 Mean: 330, std: 451 Mean: 848, std: 1396
-Locked 7 2 Mean: 1551, std:1361 Mean: 1631, std: 1837
-Table 3: Engagement statistics of 100 sampled scholars.
-In addition to segmentation, many scholars direct their followers to Threads and Bluesky instead of Mastodon, with 2% providing Threads information and 6% providing Bluesky information among the sampled scholars.
-This trend indicates that Mastodon faces significant competition from other emerging platforms, making it even more challenging for Mastodon to become a preferred alternative for academic discourse.
-In summary, we observe that the majority of scholars are either solely using Twitter or using both Twitter and Mastodon simultaneously.
-Comparing the follower and following counts across these platforms, Mastodon hosts a smaller community.
-Competition from other emerging social media platforms further limits Mastodon’s growth, indicating that scholars are exploring various alternatives rather than committing to Mastodon.
-We cross-mapped Twitter activity levels with Mastodon role categories, as shown in Table 4.
-Our analysis reveals that 55% of the sampled scholars are relatively active on both Mastodon (long-term or persistent users) and Twitter (active users).
-Among those who are inactive on Twitter or have locked their accounts, 71% are long-term or persistent users of Mastodon.
-This indicates that while Mastodon has some loyal users, their numbers are relatively small compared to the entire academic community.
-One_Time Short_Term Long_Term Persistent
-Active 12 19 28 27
-Inactive 1 1 2 1
-protected 0 2 1 6
-Table 4: Mapping of Twitter (Active, Inactive, Protected) and Mastodon (One Time, Short Term, Long Term, Persistent) activity levels.
-4 Discussion and Conclusion
-Our analysis of professional communities migrating between social media platforms highlights the unique challenges these users face during such transitions.
-While decentralized platform offered a promising alternative, the established communities and clear audience engagement on Twitter proved too significant to overcome.
-The fragmented communities on Mastodon lacked the cohesiveness found on Twitter.
-Our results show that the initial surge in Mastodon sign-ups did not translate into sustained user engagement.
-Among the users we tracked, the majority failed to maintain the same level of activity.
-This could be attributed to steep learning curve compared to twitter as well as competition from more user-friendly platforms like Bluesky and Threads.
-These alternatives might have drawn away potential users exacerbating Mastodonś difficulties in becoming a central hub for professionals communities.
-Additionally, those who continued to post consistently on Mastodon, did not receive same level of engagement as on Twitter which is crucial for sustaining professional discourse.
-Our study highlights that without a good alternative, professionals will naturally gravitate back to familiar platforms where their communities are already established through years.
+”I’m curious re categories of people I follow and where they go.
+IT people seem to have mostly gone to mastodon; librarians ditto; academic leftists stayed here longest, now going to bluesky (prefer commercial sites?); archaeologists have vanished for me completely (threads?).”
+”Hey if anyone’s still here on science Twitter and wondering where a lot of people have gone, they’re trying out bluesky and liking it. This time it’s real (not like mastodon). #academic”
+From a qualitative perspective, these conversations reflect both optimism in Mastodon’s potential and the practical concerns academics face during the migration.
+5. Discussion and Conclusion
+In this work, we explored the migratory dynamics of Academic Twitter communities to Mastodon.
+Through a series of analyses, we highlight that building a community on a new platform requires more than a coordinated effort and availability of alternatives.
+Results from longitudinal analysis of Mastodon user data showed that the initial surge in sign-ups did not translate into sustained long-term user engagement.
+Most recent user activity on Twitter reveals that many academics in our dataset returned to using Twitter actively.
+Further analysis of engagement metrics reveals that the level of established history, as well as the strong communities established on Twitter, with some over a decade, proved too significant to overcome.
+Without a complete migration of user history, and established community, professionals will naturally gravitate back to familiar platforms where their communities are already established through years.
+Several factors likely contributed to the failure of Mastodon to retain users.
+First, its decentralized structure comes with a steep learning curve compared to Twitter.
+Furthermore, competition from more user-friendly platforms like Bluesky and Threads potentially drew users away from Mastodon.
+Additionally, those who continued to post consistently on Mastodon, did not receive same level of engagement as on Twitter.
+While some users expressed doubts, certain groups of academics stayed committed to Mastodon, making this migration less of a straightforward failure and more of a complex transition.
+One significant limitation of this study is the inconsistency in data collection timing across different platforms.
+This discrepancy could potentially introduce biases in the analysis, as activity levels of scholars might vary depending on when the data was captured.
+Additionally, incorrect or incomplete social media handles provided by the scholars led to gaps in data that we could not manually resolve due to the scale of the dataset.
+Moreover, the changes to platform access (such as Twitter API rate limitations) restricted our ability to gather sufficient data throughout the study.
+However, we believe our findings from the remaining data is substantial and provides valuable insights for the cross platform analyses.
 Future efforts that support migrations to decentralized platforms must address these core issues, foster community building by reducing friction, and considering network effects to retain user engagement.
 While the idea of decentralization of social media platforms holds a significant promise, practical challenges persist that require more than just technical solutions.
-Successful transitioning involves creating communities that rival established social media networks and user experiences of platforms such as Twitter.
-{credits}
-[…]
-There is no competing interest.
+However, the Academic Twitter community’s large scale attempt to migrate from Twitter demonstrates their adaptability and resilience.
+If any group could successfully make this migration, it would be a community like Academic Twitter, united by a shared purpose and strong networks.
+Yet, as encouraging as these efforts are, there is a long way to go.

### v3

@@
-The Failed Migration of Academic Twitter
-Abstract.
-Following changes in Twitter’s ownership and subsequent changes to content moderation policies, many in academia looked to move their discourse elsewhere and migration to Mastodon was pursued by some.
-Our study looks at the dynamics of this migration.
-Utilizing publicly available user account data, we track the posting activity of academics on Mastodon over a one year period.
-We also gathered follower-followee relationships to map internal networks, finding that the subset of academics who engaged in migration were well-connected.
+The Failed Migration of Academic Twitter: A Case Study of Precocious Adopters
+Abstract
+Following changes in Twitter’s ownership in 2022 and subsequent changes to content moderation policies, many in academia looked to move their discourse elsewhere and migration to Mastodon was pursued by some.
+Our study examines the behavior of a self-organized group of early academic adopters who joined Mastodon following changes in Twitter’s ownership.
+Utilizing publicly available user account data drawn from a voluntarily curated list of academics, we track the posting activity of these early adopters on Mastodon over a one year period.
+We also study follower-followee and interaction relationships to map internal networks, finding that the subset of academics who migrated to Mastodon were well-connected.
 However, this strong internal connectivity was insufficient to prevent users from returning to Twitter/X.
-Our analyses reveal significant challenges sustaining user engagement on Mastodon due to its decentralized structure as well as competition from other platforms such as Bluesky and Threads.
-The movement lost momentum after an initial surge of enthusiasm where the main network was fully established as most users did not maintain their activity levels, and those who did faced lower levels of engagement.
-Our findings highlight the challenges involved in transitioning professional communities to decentralized platforms, emphasizing the need for focus on community building for long-term user engagement.
-Social Media, Mastodon, Twitter, Social Movement, Computational Social Science, Decentralized Platforms
-copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2024; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Social networking sitesccs: Human-centered computing Social mediaccs: Human-centered computing Computer supported cooperative workccs: Human-centered computing Social networks
-1. Introduction
+Our analyses show that early adopters struggled to maintain engagement, shaped by Mastodon’s decentralized design and competition from alternatives such as Bluesky and Threads.
+The migration effort lost momentum after an initial surge, as most early adopters reduced activity or returned to Twitter.
+Our survival analysis further reveals that retention is strongly linked to diverse cross-server engagement and topic-server communities.
+Users with large pre-existing Twitter presence face significantly higher attrition risk, highlighting the challenge of replicating established social connections in decentralized ecosystem.
+By examining the coordinated migration attempt of early adopters, we find that even this highly motivated group faced substantial challenges, suggesting that later or less coordinated efforts would likely encounter even greater barriers.
+Introduction
 Social media platforms have become essential outlets for communication, information dissemination, and professional networking.
-Among these platforms, Twitter (now known as “X”) has gained significant popularity and has been likened to a virtual town square for its role facilitating public discourse (Mascaro and Goggins, 2012).
-Academics and journalists populate sub-communities on Twitter and use the platform for professional communication (Jurkowitz and Gottfried, 2022; Malik et al., 2019).
-Academics have extensively used Twitter to share research findings, announce academic events such as conferences, and contribute to wider scholarly discourse (Mohammadi et al., 2018; Ebner, 2013; Mahrt et al., 2014).
-However, after Elon Musk’s acquisition of the company and controversial changes to content moderation policies, many on so-called Academic Twitter sought alternatives (Cava et al., 2023).
-Among academics, Mastodon, a decentralized social media platform, emerged as a popular alternative with some creating servers themed around their areas of interest.
-Unlike traditional platforms, Mastodon operates on a federated model where independent servers interact through a common protocol, allowing users to join different instances based on their preferences while still communicating across the entire network (Zignani et al., 2018).
-This structure aims to reduce centralized control over internet services by distributing data and services across multiple nodes, enhancing privacy, security, and resilience against censorship and corporate control (Raman et al., 2019).
-Yet despite these advantages, a large-scale measurement study of Mastodon revealed that this decentralized structure also leads to lower user engagement compared to centralized platforms like Twitter, where network effects and a larger, more active user base drive higher engagement levels (Raman et al., 2019).
+Among these platforms, Twitter [^1] Note on terminology: Twitter rebranded to X in July 2023, which occurred midway through our data collection.
+For simplicity, we use the terminology Twitter and tweets throughout the paper.
+(now known as “X”) has gained significant popularity and has been likened to a virtual town square for its role facilitating public discourse (Mascaro and Goggins 2012).
+Academics and journalists populate sub-communities on Twitter and use the platform for professional communication (Jurkowitz and Gottfried 2022; Malik et al.
+2019).
+In this work, we use the term Academic Twitter not to denote a formal organization, but as a shorthand for a diffuse network of scholars who use the platform for professional discourse.
+Academics have extensively used Twitter to share research findings, announce academic events such as conferences, and contribute to wider scholarly discourse (Veletsianos 2012; Mohammadi et al.
+2018; Ebner 2013; Holmberg and Thelwall 2014).
+However, after Elon Musk’s acquisition of the company and controversial changes to content moderation policies, many on so-called Academic Twitter sought alternatives (Cava et al.
+2023).
+Among academics, Mastodon emerged as a popular alternative, with communities organized around shared interests.
+Unlike traditional platforms, Mastodon operates on a federated model in which independent servers communicate through a shared protocol, allowing users to join different instances while remaining connected across the network (Zignani et al.
+2018).
+This structure aims to reduce centralized control over internet services by distributing data and services across multiple nodes, enhancing privacy, security, and resilience against censorship and corporate control (Raman et al.
+2019).
+Yet despite these advantages, this decentralized structure also leads to lower user engagement compared to centralized platforms like Twitter, where network effects and a larger, more active user base drive higher engagement levels (Raman et al.
+2019).
 The migration of users between social media platforms has been a topic of interest among researchers, particularly in the context of changes in platform policies, user experiences, and factors that influence user decisions and the implications for online communities.
-Social media movements often occur in response to significant changes in platform policies, ownership, user dissatisfaction or perceived deficiencies in a platform (Xu et al., 2014).
-Qualitative analyses, including interviews and surveys of users migrating from platforms like Fandom to alternatives such as Tumblr, reveal that policy and value-based aspects are primary catalysts of change (Fiesler and Dym, 2020).
-Additionally, research has shown that user migration is influenced by perceived platform attractiveness and peer effects (Hou and Shiau, 2020).
-A study of users migrating from Twitter to Mastodon post-Musk’s acquisition indicated that social influence, network structure, commitment, and shared identity significantly drive collective behavior changes in online communities (Cava et al., 2023).
-Analyses showed Twitter-to-Mastodon migration was characterized by low Twitter account deletion rates, concentration of users in large Mastodon instances, varied posting behaviors across platforms (He et al., 2023).
-A similar study on Twitter-to-Mastodon migration revealed lower user engagement and posting activity on Mastodon, highlighting challenges in retaining active users due to its decentralized architecture (Jeong et al., 2023).
-Despite prior work characterizing the general population of platform users, there remains a gap in targeted research concerning specific professional communities.
-In this study, we address this gap by examining the migration patterns and behaviors of academic users, providing a more nuanced understanding of how these particular communities navigate platform transitions.
-Given the significant role that Twitter plays in an academic’s professional life (Knight and Kaye, 2016; Klar et al., 2020; Stewart, 2015), academics represent a uniquely valuable group for studying the migratory dynamics.
-Our study addresses the following research questions:
-RQ1: To what extent were academics engaged in the migration from Twitter to Mastodon, and how can the success or failure of this migration be assessed?
-RQ2: Which academic fields have demonstrated the highest levels of persistence on Mastodon, and what factors characterize their migration experiences?
-Our findings indicate that majority of academic researchers who moved to Mastodon failed to maintain their posting activity and follower engagement over long term.
+Social media movements often occur in response to significant changes in platform policies, ownership, user dissatisfaction or perceived deficiencies in a platform (Xu et al.
+2014).
+Qualitative analyses, including interviews and surveys of users migrating from platforms like Fandom to alternatives such as Tumblr, reveal that policy and value-based aspects are primary catalysts of change (Fiesler and Dym 2020).
+Beyond platform-level triggers, prior studies highlight the role of social dynamics in shaping migration behavior.
+User migration has been linked to perceived platform attractiveness and peer effects (Hou and Shiau 2020).
+In the context of Twitter’s ownership change, migration to Mastodon reflected social influence, network structure, commitment, and shared identity, while remaining characterized by low Twitter account deletion rates, concentration of users in large Mastodon instances, and divergent posting behaviors across platforms (Cava et al.
+2023; He et al.
+2023).
+Recent work raises questions about sustainability, reporting lower engagement and posting activity on Mastodon compared to Twitter, and pointing to challenges in retaining active users within a decentralized architecture (Jeong et al.
+2023).
+Despite prior work characterizing general trends across the population of platform users, there remains a lack of targeted research concerning migration of narrower and highly interconnected communities characterized by shared motivations and coordinated efforts.
+We address this gap by providing a targeted case study of early academic adopters responding to platform changes, examining their migration patterns and behaviors to offer a more nuanced understanding of how these users navigate platform transitions.
+Given the significant role that Twitter plays in many academics’ professional lives (Knight and Kaye 2016; Klar et al.
+2020; Stewart 2015), academics represent a valuable subpopulation for studying migratory dynamics.
+Our work addresses the following research questions:
+RQ1: How did early academic adopters navigate the migration to Mastodon, and what do longitudinal, network, and cross-platform patterns reveal about their retention?
+RQ2: Which measurable user and community characteristics are associated with long-term retention on Mastodon?
+Our findings indicate that the majority of academic researchers who moved to Mastodon failed to maintain their posting activity and follower engagement over the long term.
 Although migrated academic accounts formed strongly-connected internal networks, these connections were insufficient to sustain user engagement over time.
 Of note, however, communities that established field specific server instances had higher user retention rate compared to those on general-purpose servers.
-Cross platform analysis of most recent Twitter activity showed that most users returned to Twitter.
-These results underscore the challenges involved and the potential opportunities to build and foster sustainable communities at scale on new platforms.
-2. Related Work
-2.1. Characteristics of Mastodon
-The decentralized structure Mastodon and similar platforms leads to unique user engagement dynamics that are different from centralized platforms.
-Research suggests that users are drawn to Mastodon for the ease of protecting their information from data mining and the absence of algorithms shaping their feeds (Lee and Wang, 2023).
-However, this can also make content discovery more difficult (Zignani et al., 2019).
-The decentralized structure encourages strong, tight-knit communities within individual server instances but fewer connections across instances (Zignani et al., 2018).
+Cross-platform analysis of most recent Twitter activity showed that most users returned to Twitter.
+Survival analysis confirms that user’s persistence is associated with a number of factors including diverse cross-server engagement, early platform activity, choice of server.
+Furthermore, cross-platform modeling shows a divergent pattern with long history of prior activity on Twitter is associated with higher attrition risks.
+These results underscore the challenges and opportunities associated with building and maintaining communities on new platforms.
+Related Work
+Cross-Platform Migration
+Social media movements often occur in response to significant changes in platform policies, ownership, user dissatisfaction or perceived deficiencies in a platform (Xu et al.
+2014).
+Early work analyzed it using Push Pull Moore framework with Push effects (e.g., policy changes) drive users away, while Pull effects (e.g., better features) attract them to alternatives.
+However, Mooring factors such as switching costs often prevent complete migration (Chang et al.
+2014; Lin and Wang 2017).
+Qualitative analyses, including interviews and surveys of users migrating from platforms like Fandom to alternatives such as Tumblr, reveal that policy and value-based aspects are primary catalysts in the migration process (Fiesler and Dym 2020).
+Additionally, research has shown that user migration is influenced by factors such as perceived platform attractiveness and peer influence (Hou and Shiau 2020).
+Study of users migrating from Twitter to Mastodon post-Elon Musk’s acquisition indicated that social influence, network structure, commitment, and shared identity significantly drive collective behavior changes in online communities (Cava et al.
+2023).
+Analysis showed Twitter to Mastodon migration is characterized by low Twitter account deletion rates, concentration of users in large Mastodon instances, varied posting behaviors across platforms (He et al.
+2023).
+A similar study on Twitter-to-Mastodon migration reveals lower user engagement and posting activity on Mastodon, highlighting challenges in retaining active users due to its decentralized architecture (Jeong et al.
+2023).
+These results are consistent with evidence that the transition is often an incomplete with users maintaining cross platform presence rather than fully substituting one platform to the other (Peña-Fernández et al.
+2025).
+More broadly recent work on platform switching from Twitter shows that users disperse across multiple destination platforms rather than relocating to a single successor platform (Radivojevic et al.
+2025).
+Our work looks more deeply at the specific case of the relatively well-synchronized and coordinated movement of communities of academic users across platforms.
+Deplatforming as Involuntary Migration
+Unlike voluntary platform migration, which is often motivated by user dissatisfaction or policy changes, deplatforming represents a form of forced migration where users are compelled to leave due to removal or suspension (Rogers 2020).
+A study of Reddit’s 2015 ban shows that deplatforming can reduce the visibility of harmful content on mainstream platforms while simultaneously displacing communities to alternative or less moderated spaces (Chandrasekharan et al.
+2017).
+Parler was deplatformed after the 2021 U.S.
+Capitol attack, while Gab expanded as a refuge for banned Twitter users; however, research shows that such platforms often attract smaller or more fragmented audiences and face challenges related to moderation and community stability (Aliapoulios et al.
+2021; Zannettou et al.
+2018).
+While distinct from voluntary migration, deplatforming likewise underscores how platform-level conditions shape user mobility and contextualizes the broader challenges communities face when establishing presence on alternative platforms.
+Characteristics of Mastodon
+The decentralized structure of Mastodon and similar platforms leads to unique user engagement dynamics that are different from centralized platforms.
+Research suggests that users are drawn to Mastodon for the ease of protecting their information from data mining and the absence of algorithms shaping their feeds (Lee and Wang 2023).
+However, this can also make content discovery more difficult (Zignani et al.
+2019) and may contribute to network fragmentation (Kleppmann et al.
+2024), shaping interactions and communication patterns that differ from centralized platforms (Brauweiler et al.
+2025).
+The decentralized structure encourages strong, tight-knit communities within individual server instances but fewer connections across instances (Zignani et al.
+2018).
 These characteristics pose challenges for migration as well.
-2.2. Cross-platform migration
-Social media movements often occur in response to significant changes in platform policies, ownership, user dissatisfaction or perceived deficiencies in a platform (Xu et al., 2014).
-The migration of users between social media platforms has been a topic of interest among researchers, particularly in the context of changes in platform policies, user experiences, and factors that influence user decisions and the implications for online communities’ retention.
-Qualitative analyses, including interviews and surveys of users migrating from platforms like Fandom to alternatives such as Tumblr, reveal that policy and value-based aspects are primary catalysts in the migration process (Fiesler and Dym, 2020).
-Additionally, research has shown that user migration is influenced by factors such as perceived platform attractiveness and peer influence (Hou and Shiau, 2020).
-Study of users migrating from Twitter to Mastodon post-Elon Musk’s acquisition indicated that social influence, network structure, commitment, and shared identity significantly drive collective behavior changes in online communities (Cava et al., 2023).
-Analysis showed Twitter to Mastodon migration is characterized by low Twitter account deletion rates, concentration of users in large Mastodon instances, varied posting behaviors across platforms (He et al., 2023).
-A similar study on Twitter-to-Mastodon migration reveals lower user engagement and posting activity on Mastodon, highlighting challenges in retaining active users due to its decentralized architecture (Jeong et al., 2023).
-Our work looks more deeply at the specific case of the relatively well-synchronized and coordinated movement of communities of academic users across platforms.
-3. Data
-Note on terminology: Twitter rebranded to X in July 2023, which occurred midway through our data collection. For simplicity, we use the terminology Twitter and tweets throughout the paper.
+Our study examines how specialized communities—particularly academics—navigated these dynamics in their movement from Twitter to Mastodon.
+Data
 Our analyses are primarily based on a public Github repository containing lists of academics on Mastodon [^2] https://github.com/nathanlesage/academics-on-mastodon.
@@
 Some lists also included each user’s Twitter handle.
-While the database continues to expand, for the purpose of longitudinal analysis, we focused only on scholars whose Mastodon accounts were already in the repository at the beginning of November 2022, noting the Twitter-to-Mastodon migration began in late October 2022.
+For the purpose of longitudinal analysis, we focused only on scholars whose Mastodon accounts were already in the repository at the beginning of November 2022, noting the Twitter-to-Mastodon migration began in late October 2022.
 That is, the users in our dataset were part of a coordinated, synchronized migration effort in November 2022.
 In total, we collected the Mastodon profiles and corresponding public metrics, e.g., number of followers, following, timestamps of most recent status of 7,542 academic user accounts using a customized Mastodon API [^3] https://github.com/halcy/Mastodon.py.git.
-Data was collected weekly from November 2022 to October 2023 allowing for temporal analysis of user activity and engagement over the span of a year.
+We collected data weekly from November 2022 to October 2023 allowing for temporal analysis of user activity and engagement over the span of a year.
 In addition, we collected the most recent Mastodon data related to follower and following accounts for each scholar to conduct network analyses based on fields of study and associated servers.
+Additionally, to capture dynamics beyond follower and followee connections, we collected the user interactions (replies, boosts, mentions, and favorites) covering the full observation window from November 2022 to October 2023.
+We collected activity for users in our sample to extract both outgoing interactions and incoming interactions.
+Note that this interaction dataset was collected retrospectively in December 2025.
 For cross-platform analyses, we used self-reported Twitter handles to gather scholars’ profile information.
@@
 Additionally, we used the Twitter ID collection endpoint to extract the posting time of their most recent tweet.
-Finally, to understand discussions surrounding Mastodon in academic context, we collected tweets during the one year window via Zeeschuimer [^3] https://github.com/digitalmethodsinitiative/zeeschuimer using the boolean search “mastodon (researcher OR researchers OR academic OR academics OR professor OR professors OR scholar OR scholars) lang:en until:2023-11-01 since:2022-10-01”.
-Since many tweets were included because the username contained the word “mastodon” (as some users use their Mastodon account name as their Twitter username) we further filtered out these tweets.
+Finally, to understand discussions surrounding Twitter migration to Mastodon on both platforms, we collected posts from both platforms during the one year window.
+Specifically, we collected tweets via Zeeschuimer [^4] https://github.com/digitalmethodsinitiative/zeeschuimer using the boolean search “mastodon (researcher OR researchers OR academic OR academics OR professor OR professors OR scholar OR scholars) lang:en until:2023-11-01 since:2022-10-01”.
+The set of academic-related terms was developed through a preliminary scan of tweets, ensuring coverage of both individual (e.g., “professor,” “scholar”) and collective (e.g., “academics,” “researchers”) identifiers.
+Since many tweets were included because the display name contained the word “mastodon” (as some users use their Mastodon account name as their Twitter display name) we further filtered out these tweets.
 We retained only those that contained relevant keywords in the main body of the tweet, resulting in a total of 1,497 tweets.
-Figure 1. Waterfall plot showing number of monthly active users (active=at least one post per month).
-4. Results
+We also collected posts through the Mastodon API using migration-related hashtags adapted from prior work (Jeong et al.
+2024).
+We excluded the tag ”MastodonSocial” because it is commonly used for general content on Mastodon rather than migration.
+From a total pool of 12,654 migration related posts, we then applied a second stage filter using the same set of academic keywords described earlier to identify content relevant to academic migration.
+This process yielded 129 posts in the final dataset.
+RQ1: Migration Dynamics and User Retention
 Following, we present findings from multiple analyses conducted to understand the dynamics of the migration.
@@
 Then, we used network analysis to understand the connectivity and distribution of academics across Mastodon servers.
-Next, we did a cross-platform analysis using the subset of users who provided Twitter handles comparing their activity patterns on Twitter and Mastodon.
-Finally, we ran exploratory analyses on Twitter posts related to the migration to understand user attitudes towards migration.
-4.1. Longitudinal Analysis
-4.1.1. Attrition of academic accounts on Mastodon
-Figure 1 gives a count waterfall plot illustrating the change in number of active users aggregated by month and a line plot showing the monthly retention rate based on the previous month.
+Next, we conducted a cross-platform analysis using the subset of users who provided Twitter handles comparing their activity patterns on Twitter and Mastodon.
+Taken together, these analyses identify patterns and potential factors associated with user retention, which in turn motivate the selection of features included in the survival analysis for RQ2.
+Longitudinal Patterns of Adoption and Retention
+Attrition of academic accounts on Mastodon
+Figure 1 (Top) gives a bar plot illustrating the change in number of active users aggregated by month and a line plot showing the monthly retention rate based on the previous month.
 In this context, we considered a user to be active if they made at least a post in a given month.
@@
 [^5] There were 1, 4, and 32 scholars had their last activity in April, May, and October 2022, respectively, which fell outside our collection window.
-These scholars were excluded from the waterfall plot but included in subsequent analysis.
+These scholars were excluded from the bar plot but included in subsequent analysis.
 Over the next seven months, the number of active scholars experienced a decline, showing a minor resurgence in July 2023 before continuing its descent, ultimately reaching 2,398 by October 2023.
-We observe a subtantial rate of attrition; each month, approximately 10% to 20% of scholars in our set stopped actively using the platform.
-4.1.2. Engagement with academic accounts on Mastodon
-Given the central role of engagement in user activity, we look at the engagement with migrated academic accounts on Mastodon. To do so, we first characterize accounts based on their activity.
+We observe a substantial rate of attrition; each month, approximately 10% to 20% of scholars in our set stopped actively using the platform.
+The surge in July 2023 coincided with significant shifts in Twitter’s policy on limiting number of posts user can view and officially rebranded to “X” which may have prompted a renewed but temporary interest in Mastodon as alternative.
+Engagement with academic accounts on Mastodon
+Given the central role of engagement in user activity, we look at the engagement with migrated academic accounts on Mastodon.
+Figure 1 (Bottom) presents the distribution of user active spans, which we use to characterize accounts into activity patterns for comparison across dimensions in later sections.
+Figure 1: Top: Histogram showing number of monthly active users (active=at least one post per month).
+Bottom: Histogram showing distribution of user active spans (number of days they are on Mastodon aggregated by week).
 Active users.
@@
 For our purposes, we’ve defined this initial period between October 15 and November 30, 2022.
-We further divide active users into sub-categories based on level of engagement: one-time visitors; short-term, long-term, and persistent adopters.
-Each category reflects the duration of their activity on Mastodon following the migration.
+We categorize active users into four engagement groups based on the empirical distribution of active spans using the 25th, 50th, and 75th percentiles: one-time visitors; short-term, long-term, and persistent adopters.
+These data-driven cutoffs reflect the underlying heterogeneity in the duration of their activity on Mastodon following the migration.
+The days cutoffs are rounded to the nearest integer.
 One-time visitors.
-These academics created an account on Mastodon at the beginning of the Twitter-to-Mastodon migration, posted only within the first week of joining and then abandoned the platform ($n_{days}\leq 7$).
-This brief engagement suggests a possible exploration of the platform following the migration advocacy, without finding enough value or community to continue their activity.
+These academics engaged only briefly at the beginning of the Twitter-to-Mastodon migration, with active spans falling below the 25th percentile of the overall distribution ($n_{days}\leq 39$).
+Their minimal participation suggests exploratory behavior during the migration wave without continued investment in the platform.
 Short-term adopters.
-This group includes those who actively used Mastodon for more than one week but did not persist beyond 100 days.
-They left the platform possibly due to unmet expectations or a stronger pull back to more familiar platforms like Twitter.
-Long-term adopters. These users remained active on Mastodon from the time of migration for relatively long time ($100<n_{days}\leq 300$) but have been inactive.
+This group falls between the 25th and 50th percentiles of active spans ($39<n_{days}\leq 222$).
+These users engaged beyond initial exploration but did not sustain participation long enough to become stable contributors, potentially reflecting unmet expectations or competing commitments on other platforms.
+Long-term adopters.
+These users fall between the 50th and 75th percentiles of the distribution ($222<n_{days}\leq 329$).
+They demonstrated extended engagement following the migration but did not remain active through the full analytic period.
 Persistent adopters.
-These users remained active on Mastodon from initial migration though the end of our analyses ($n_{days}>300$).
-Their sustained engagement indicates successful integration into the Mastodon community.
-4.1.3. Inactive users.
+These users fall above the 75th percentile ($n_{days}>329$), corresponding to the longest active spans through the end of our analyses.
+Their sustained engagement suggests successful integration into Mastodon’s academic community and continued participation over time.
+Inactive users.
 Academics in this category were already using Mastodon before the migration wave from Twitter.
@@
 The distribution of users is given in Table 1.
-Note that only 35.4% of academic users in our dataset are active and consistent in their use of Mastodon.
-The majority demonstrate sporadic engagement.
 User Category Subcategory Count
-Active One-time visitors 1,183
-Short-term adopters 1,624
-Long-term adopters 2,065
-Persistent adopters 2,665
+Active One-time visitors 1895
+Short-term adopters 1876
+Long-term adopters 1921
+Persistent adopters 1845
 Inactive One-Time Visitors 4
 Returned Users 1
-Table 1. Number of users, by activity.
-We calculate the average following, follower and post counts for each user category, aggregated by month show in Figure 2.
+Table 1: Number of users, by activity.
+We calculate the average following, follower and post counts for each user category, aggregated by month as shown in Figure A1.
 Our analysis reveals a sharp increase in all three public metrics at the onset of the migration period, specifically from November 2022 to December 2022.
@@
 The persistent users’ continued efforts to engage and post suggest a commitment to the platform, yet the limited increase in followers indicates potential barriers in user retention and community building.
-Figure 2. Average public metrics and statuses count for active users.
-4.1.4. Field-specific activity differences among academic accounts on Mastodon
-As discussed in Section 3, users self-identified their research disciplines in public lists.
-We aggregate these disciplines into 9 broad fields, consistent with existing taxonomies of academic disciplines (bepress., 2019).
+Field-specific activity differences among academic accounts on Mastodon
+As discussed in Section Data, users self-identified their research disciplines in public lists.
+We aggregate these disciplines into 9 broad fields, consistent with existing taxonomies of academic disciplines (bepress.
+2019).
+We adopt this taxonomy because its granular disciplinary structure aligns more closely with the self-reported research areas in our dataset, and it is widely deployed in institutional repository metadata system (e.g.,  [^6] https://guides.lib.odu.edu/c.php?g=502940&p=5839940, [^7] https://stars.library.ucf.edu/faq.html, [^8] https://digitalcommons.lib.uconn.edu/about.html).
 Scholars identifying with multiple disciplines and open science researchers were categorized as multi-disciplinary researchers.
@@
 Disciplinary statistics of the 7,537 academics in our datasets are provided in Table A1 in the Appendix.
-We identify the proportion of one-time, short-term, long-term, and persistent users by field (see Table 2).
-The highest proportion of persistent users is observed within multidisciplinary researchers.
-The public metrics aggregated by field and by month are shown in Figure A1 in the Appendix.
-Researchers in the physical sciences and mathematics have notably higher average number of posts, indicating their consistent and significant contribution to discussions on Mastodon.
-Scholars in medicine and health sciences exhibit relatively less frequent posting.
-Field One Time (%) Short Term (%) Long Term (%) Persistent (%) Total Percent In-Degree Out-Degree
-Social & Behavioral Sciences 15.7 21.4 27.9 35.0 2918 38.72 98881 91958
-Arts & Humanities 20.5 19.4 24.8 35.3 1522 20.19 38692 39986
-Physical Sciences & Mathematics 12.0 21.2 27.8 39.0 851 11.29 11453 11980
-Life Sciences 10.7 20.2 41.7 27.4 84 1.11 338 648
-Education 21.3 26.2 24.2 28.3 244 3.24 2249 4171
-Law 17.9 28.3 26.6 27.3 587 7.79 21664 27450
-Medicine & Health Sciences 10.8 30.3 29.3 29.6 287 3.81 4292 4923
-Multidisciplinary 10.8 18.3 28.8 42.1 1040 13.80 56371 52818
-Business 25 25 25 25 4 0.05 60 66
-Table 2. Distribution of user engagement and node degree statistics by academic field.
-4.2. Mastodon Network Analyses
-To understand the connectivity within and between different fields, we collected the complete follower-followee data of all academic accounts on Mastodon, and constructed an internal connection network within each academic group, as well as a network aggregated by field of study.
-4.2.1. Internal Network Aggregated by Field
+We identify the proportion of one-time, short-term, long-term, and persistent users by field (see Table A2 in the Appendix).
+The highest proportion of persistent users is observed among multidisciplinary researchers.
+One possible explanation is that interdisciplinary scholars engage with discussions spanning multiple fields, which may provide a broader set of conversational opportunities and connections that support continued participation.
+This observation motivates the inclusion of a binary indicator for interdisciplinary research in our subsequent modeling of user retention.
+Network Structure and Community Formation
+To understand the connectivity within and between different fields, we collected the complete follower-followee network and multiple types of interaction networks of all academic accounts on Mastodon, and constructed an internal connection network within each academic group, as well as a network aggregated by field of study.
+Internal networks aggregated by field
 Figure 2 shows the network connections across different fields.
@@
 Our analysis reveals that the academic migration has led to the formation of a well-connected internal network, suggesting that researchers actively attempted to establish a new community on Mastodon.
-However, despite these coordinated efforts, the migration still faced significant challenges, as reflected through the low retention rates shown in Section 4.1.
-4.2.2. Internal Networks aggregated by Server
+However, despite these coordinated efforts, the migration still faced significant challenges, as reflected through the low retention rates shown in Section Longitudinal Analysis.
+Figure 2: Internal field network (size: number of scholars in each field; edge weights: frequency of connections; edge color: same as the source; self-loops are omitted for better visualization.)
+Figure 3: (Top) Server network constructed from internal follower–followee relationships.
+Isolated servers are removed for clarity.
+Node size is scaled by the number of scholars within each server.
+Servers with more than 100 scholars are highlighted in red.
+(Bottom) Internal comment interaction network among servers with more than 100 scholars.
+Node size is scaled by the total number of interactions for each server, and edge width is scaled by the number of interactions between servers.
+Internal boost, favorite, and mention networks are reported in Figure A2 in the Appendix.
+Internal networks aggregated by server
 Prior studies have shown that the Mastodon’s server instance based structure makes it easier for users to discover and connect with others within the same instance than outside instance.
-This leads to stronger within instance ties compared to weaker links across instances (Zignani et al., 2019).
-Therefore, we analyzed the follower-followee network aggregated by Mastodon instances, shown in Figure 4.
+This leads to stronger within instance ties compared to weaker links across instances (Zignani et al.
+2019).
+Therefore, we analyzed the follower-followee network aggregated by Mastodon instances and the corresponding interaction networks among top servers, shown in Figure 3.
 The academic users were spread across a total of 342 servers, with 9 server instances being isolates without any connecting edges to other instances, which were removed for clarity.
@@
 During the migration, different disciplines used different strategies: some used general-purpose servers like mastodon.social which host a wide variety of content beyond just academic content, while others took a coordinated approach on setting up their field specific server instances such as infosec.exchange and journa.host.
-Figure 4 shows the follower-followee connections across different servers showing certain general purpose instances such as mastodon.social are popular among the tracked academic users (containing 21.3% of the total) than other general purpose servers like mas.to.
-Figure 3. internal field network (size: number of scholars in each field; edge weights: frequency of connections; edge color: same as the source).
-4.2.3. Case study: Mastodon Community of Information Security
-We also examined a discipline-specific server, unique to decentralized social platforms, and its user retention rate using a case study of the Information Security discipline.
-We chose this discipline based on our field-level engagement analysis shown in Figure A1 in Appendix, in the Appendix, which indicates that scholars in physical sciences and mathematics were among the most active on Mastodon.
-To further understand the impact of server instance choice on community formation, we visualized their internal network, highlighting the top five servers where these scholars are based.
-Figure A2 shows the internal follower-followee network for physical science and mathematics field, highlighting the top five servers from a total of 33 servers that these scholars use.
-Interestingly, scholars from Information Security formed a strongly connected community on a custom discipline specific server, unlike other disciplines that used general servers such as mastodon.social or fediscience.org.
-To explore how a discipline specific server could influence user retention and engagement, we isolated the internal network of accounts within Information Security discipline.
-Figure 5 shows the internal network of the discipline, where most scholars congregated on a single server: infosec.exchange (highlighted in red).
-We compare the retention roles of scholars from the discipline with those from a discipline where users were concentrated in a general-purpose server and other sub-fields of similar size.
-We see a significantly high percentage of persistent users for information security scholars (56.8%) in comparison to other comparable disciplines.
+Figure 3 (Top) shows the follower-followee connections across different servers showing certain general purpose instances such as mastodon.social are popular among the tracked academic users (containing 21.3% of the total) than other general purpose servers like mas.to.
+We observe a central hub pattern in which the majority of follower–followee connections involve the mastodon.social server and other servers, forming a radiating structure.
+To examine interactions beyond follower relationships, we further plot interaction networks among top servers with more than 100 scholars (highlighted in red in Figure 3 (Top)).
+Figure 3 (Bottom) shows the internal comment interaction network (boost, favorite, and mention networks are presented in Figure A2 in the Appendix).
+Node size is scaled by the total number of interactions for each server, and edge width is scaled by the number of interactions between server pairs.
+We find that even when servers are not directly connected through follower–followee relationships, they are connected through interactions such as comments, mentions, and boosts.
+This motivates the inclusion of both within-server and inter-server interaction in subsequent modeling, as these interaction-based connections capture federated engagement that is not visible in the follower network alone.
+In particular, interaction networks reveal cross-server communication pathways that operate independently of stable follower ties.
+Comparison Across Communities.
+To explore how a discipline specific server could influence user retention and engagement, we isolated the internal network of accounts within Information Security.
+Figure A3 shows the internal network of the discipline, where most scholars congregated on a single server: infosec.exchange (highlighted in red).
+We compare the retention roles of scholars from the discipline with those from a discipline where users were concentrated in a general-purpose server and other sub-fields of similar size (Education and Communication & Media Studies), as shown in Figure 4.
+We see a significantly higher percentage of persistent users amongst information security scholars (40.9%) in comparison to other comparable disciplines.
 This suggests that establishing field- or discipline-specific communities or servers could foster a stronger sense of belonging, which may help retain scholars who have migrated, reducing the likelihood of their return to Twitter or seeking alternative platforms.
-Figure 4. internal server network (removed isolated servers, red are servers with more than 100 scholars).
-Figure 5. internal follower-followee network for Information Security community (removed isolates).
-Category Info. Sec. Education Communication & Media Studies
-Count Percent Count Percent Count Percent
-One-Time 23 8.7 52 21.3 41 15.5
-Short-Term 31 11.7 64 26.2 53 20.0
-Long-Term 60 22.7 59 24.2 81 30.6
-Persistent 150 56.8 69 28.3 90 34.0
-Total 264 100.0 244 100.0 265 100.0
-Table 3. User retention breakdown for Information Security scholars compared to the sub-fields that are similar in size.
-4.3. Twitter-Mastodon Cross-platform Analysis
-Since our attrition analysis showed that only a few users remained active on Mastodon, we conducted a cross-platform analysis to analyze their account state on Twitter.
+Figure 4: User retention breakdown for Information Security scholars compared to the sub-fields that are similar in size (Info. Sec.: n=264; Comm. & Media Studies: n=265; Education: n=244.)
+Cross-Platform Patterns and Usage Shifts
+Since our attrition analysis categorized 1845 users as active on Mastodon, we conducted a cross-platform analysis to analyze their account state on Twitter.
 Approximately 50% of the scholars voluntarily provided their Twitter usernames, which were accurate and traceable to their Twitter profile pages.
@@
 Next, we retrieved the most recent posting times of these Twitter accounts to assess their activity status.
+Count Following Count Follower Count
+Active 2496 $\mu$:1989.07, $\sigma$:3087.68 $\mu$:13208.76 , $\sigma$:108495.54
+Semi-active 220 $\mu$:1137.56 , $\sigma$:1263.87 $\mu$:4467.28, $\sigma$:19152.90
+Inactive 91 $\mu$:814.56 , $\sigma$:1062.39 $\mu$:2401.46 , $\sigma$:5203.50
+Locked 324 $\mu$:1859.72 , $\sigma$:2300.91 $\mu$:6347.10 , $\sigma$:31583.98
+Table 2: Engagement statistics of 3,131 scholars (number of users per category with mean and standard deviation of following and follower counts).
 Based on their posting behavior, we categorized the Twitter account activity into the following statuses:
-Locked: Scholar’s account is locked down, meaning their posts are private and cannot be accessed.
-Inactive: Scholar’s latest tweet was posted before December 31st, 2022, and their posts were made before this date, indicating no recent activity since the start of the year.
-Semi-active: Scholar’s latest tweet was posted between December 31st, 2022 and November 1, 2023.
-This indicates some activity during the year but no recent posts after November, 2023, the last timestamp of our Mastodon tracking.
-Active: Scholar’s latest tweet was posted after November 1st, 2023, indicating continuous activity and engagement on the platform.
-Using the above categorization, we calculated the count and connection statistics for each category, and all statistics are presented in Table 4 (left).
+Locked. Scholar’s account is locked down, meaning their posts are private and cannot be accessed.
+Inactive. Scholar’s latest tweet was posted before December 31st, 2022, and their posts were made before this date, indicating no recent activity since the start of the year.
+Semi-active.
+Scholar’s latest tweet was posted between December 31st, 2022 and November 1, 2023.
+This indicates some activity during the year but no recent posts after November, 2023, the end of our Mastodon tracking.
+Active. Scholar’s latest tweet was posted after November 1st, 2023, indicating continuous activity and engagement on the platform.
+Using the above categorization, we calculated the count and connection statistics for each category, and all statistics are presented in Table 2.
 We observe that 79.72% of scholars are still active on Twitter, indicating no signs of migration or intention to fully migrate.
-These active Twitter users have accumulated a much higher number of followers, which make it more difficult to give up the platform.
+These active Twitter users have accumulated approximately 10 times more followers, which make it more difficult to give up the platform.
 Some active users still make their Mastodon information accessible to the community.
 For instance, one scholar noted:
-”Hi ya’ll for those moving over to #Mastodon you can currently find me ¡Scholar’s Mastodon Account¿.”
+“Hi ya’ll for those moving over to #Mastodon you can currently find me <<Scholar’s Mastodon Account>>.”
 This group of scholars does not personally intend to migrate but aims to maintain connections with those who do intend to fully migrate to Mastodon.
@@
 In contrast, 2.91% of the scholars still have accessible but inactive accounts since November 2022, indicating they will not return to Twitter. For example, one scholar’s last post reads,
-”I’m leaving Twitter but I’ve been advised not to delete my account in order to avoid possible misuse - so I’ll leave it here as a placeholder. See you at Mastodon!”
+“I’m leaving Twitter but I’ve been advised not to delete my account in order to avoid possible misuse - so I’ll leave it here as a placeholder. See you at Mastodon!”
 Approximately 10.35% of the scholars chose to lock their Twitter accounts, often signaling their departure from the platform by leaving a message in their profile directing followers to their Mastodon accounts.
-Figure 6. Connection metrics mapping between mastodon and Twitter.
-Count Following Count Follower Count One-time Short-term Long-term Persistent
-Active 2496 Mean:1989.07, std:3087.68 Mean:13208.76 , std:108495.54 374 578 684 860
-Semi-active 220 Mean:1137.56 , std:1263.87 Mean:4467.28, std:19152.90 39 42 43 96
-Inactive 91 Mean:814.56 , std:1062.39 Mean:2401.46 , std:5203.50 12 19 21 39
-Locked 324 Mean:1859.72 , std:2300.91 Mean:6347.10 , std:31583.98 51 51 82 140
-Table 4. Engagement statistics of 3,131 scholars and mapping of Twitter (Active, Inactive, Protected) and Mastodon (One Time, Short Term, Long Term, Persistent) activity levels.
-We cross-mapped Twitter activity levels with Mastodon role categories, as shown in Table 4 (right).
-Our analysis reveals that 49.31% of the sampled scholars are relatively active on both Mastodon (long-term or persistent users) and Twitter (active users).
-Among those who are inactive on Twitter or have locked their accounts, 67.95% are long-term or persistent users of Mastodon.
-However, this group of dedicated Mastodon users represents only 9.02% of the total scholars included in the cross-platform analysis.
-This suggests that while there is a core group of loyal users who have fully migrated to Mastodon, they make up a relatively small portion of the overall academic community.
-We also plotted the pairwise follower and following counts for the Mastodon and Twitter accounts of the same 3,131 scholars, as shown in Figure 6.
-The difference in connection scale is clear: while scholars who migrated to Mastodon made efforts to establish a new community, it is nowhere near the scale of connections they built over the years on Twitter.
+Figure 5: Cross-platform engagement mapping of Twitter and Mastodon; counts of scholars classified by Twitter status (rows) and Mastodon usage (columns).
+We cross-mapped Twitter activity levels with Mastodon role categories, as shown in Figure 5.
+Our analysis reveals that 39.95% of the sampled scholars are relatively active on both Mastodon (long-term or persistent users) and Twitter (active users).
+Among those who are inactive on Twitter or have locked their accounts, 56.14% are long-term or persistent users of Mastodon.
+However, this group of dedicated Mastodon users represents only 7.44% of the total scholars included in the cross-platform analysis.
+This suggests that while there is a core group of loyal users who have fully migrated to Mastodon, they make up a relatively small portion of the broader group of early academic adopters.
+We also plotted the follower and following counts for the Mastodon and Twitter accounts of the same 3,131 scholars and the absolute differences in the metrics, as shown in Figure A4.
+The difference in connection scale is clear: while scholars who migrated to Mastodon made efforts to establish a new community, these networks remain far smaller than the connections many of them had built on Twitter over the years.
 Additionally, we plotted the Twitter account creation date against the total number of tweets, displayed in Figure A5.
@@
 This entrenched presence highlights the difficulty of transitioning to a new platform like Mastodon without sacrificing years of accumulated influence and community ties.
-4.3.1. Alternatives to Mastodon
+In summary, we observe that the majority of scholars are either solely using Twitter or using both Twitter and Mastodon simultaneously.
+Comparing the follower and following counts across these platforms, Mastodon hosts a smaller community.
+Competition from other emerging social media platforms further limits Mastodon’s growth, indicating that scholars are exploring various alternatives rather than committing to Mastodon.
+RQ2: Modeling Long-Term Retention
+In previous sections, we have identified multiple factors that could potentially influence persistence on Mastodon.
+To understand the effect of these factors on persistence of Mastodon usage, we model them using a Cox proportional hazards model (Cox 1972).
+We selected this model to account for right censored data (users remaining active beyond the observation window) and to estimate the hazard ratios of covariates without assuming baseline hazard distribution.
+Mastodon Survival Analysis
+We used a Mastodon-centric model as our primary analysis and a cross-platform extension to understand the roles of Mastodon- and Twitter-derived features.
+The covariates used can be grouped as follows:
+Mastodon activity metrics. These represent users’ platform activity level upon migration and social network metrics: the initial number of followers, following, and posts.
+Federated interaction diversity.
+To capture how broadly a user engages across the Fediverse beyond their home server, we compute a network diversity metric from interactions.
+For each user, we extract the set of outgoing interactions and map each interaction to a target server.
+We then compute Shannon entropy (Lin 2002) over the distribution of target servers, yielding higher values when a user’s outgoing interactions are spread across many servers rather than concentrated within a single server.
+Our primary model uses diversity measured over the first 30 days of observed activity and we report sensitivity checks using alternative definitions (14 days, 60 days, and the first 50 outgoing interactions) in Figure A6 and A7.
+Account and server-level features.
+We capture users’ profile settings, the network structure of their server, and the server type.
+Specifically, we consider whether the account is locked (binary), indicating if their posts are private and/or discoverable (binary), which indicates whether a user opted into being indexed to appear in search.
+At the server level, we identify the server type (binary) as either field-/topic-specific or general purpose by manually examining their public descriptions.
+A complete list of servers is provided in Appendix A4.
+At the server level, we define in-degree ratio as the proportion of inbound connections relative to total connections and out-degree ratio as the proportion of outbound connections relative to total connections.
+Additionally, we categorize users based on their academic discipline, creating a binary indicator for Multidisciplinary scholars in Table A1 to test if these communities show a distinct retention patterns.
+Cross-platform metrics.
+These capture users’ premigration history on Twitter.
+We consider the followers and following counts on Twitter, the total number of Tweets since account creation, and the age of their Twitter account in days.
+In both models, for each account, the number of days between our first data collection (11-23-2022) and the account’s last recorded posting activity within the study window was used as survival time.
+We treat a user having experienced the event of becoming inactive if their last posting occurs before the end of our data collection (10-27-2023).
+The counts covariates are log transformed and z-score normalized to handle the covariate distribution and comparable coefficients.
+We used a penalization ($L_{2}=0.1$) to prevent overfitting.
+Mastodon-centric model
+Our first model used the complete dataset of user accounts and features extracted exclusively from Mastodon data.
+We filtered out accounts on isolated servers and those with incomplete data, reducing our sample from 7,542 to a final account count of 7,357 users.
+The model achieved a strong predictive capability (concordance of 0.69) with several factors significantly predicting the risk of becoming inactive on Mastodon.
+The initial number of a user’s followers ($HR=0.72,p<0.0005$) and number of posts ($HR=0.74,p<0.001$) were strongest protective factors; one standard deviation increase in either was associated with around 30% lower likelihood of becoming inactive (refer to Appendix Figure A6).
+The initial number of accounts followed by a user did not have a significant effect ($HR=1.02,p=0.262$).
+For the server type and profile settings, being on a topic-specific server was a strong protective factor ($HR=0.84,p<0.0005$), aligning with our observations from the information security server that tailored communities are associated with higher user retention.
+Notably, users identified as Multidisciplinary scholars exhibited higher retention rates with a hazard ratio of $0.59$ $(p<0.001)$.
+Furthermore, having a discoverable profile was also significantly linked to lower attrition ($HR=0.93,p=0.009$), suggesting that being able to be found and connected is linked to persistent usage.
+In contrast, having the profile locked was not found to be a statistically significant predictor.
+The out-degree ratio was a risk factor ($HR=1.80,p<0.001$) suggesting servers that are heavily outward looking are associated with lower user persistence.
+However, the proportion of inbound connections was not a significant predictor.
+Federated interaction diversity is a robust protective predictor.
+In the 30-day diversity model, network diversity is associated with lower attrition ($HR=0.84,p<0.001$), indicating that a one standard deviation increase in early cross-server interaction diversity corresponds to an approximately 16% lower hazard of becoming inactive.
+Sensitivity checks using alternative windows (14 days, 60 days, and first 50 outgoing interactions) yield consistent results (Appendix Table X), with effect sizes ranging from $HR=0.89$ (14 days) to $HR=0.75$ (first 50 outgoing interactions).
+Cross-platform model
+Our second model was fitted using the subset of 3,131 users for whom we had complete cross-platform features from both Mastodon and Twitter.
+This allowed us to understand how a user’s Twitter activity related to their retention on Mastodon.
+The cross-platform model achieved concordance of 0.69 indicating that addition of Twitter features does not add much overall predictive power.
+The addition of Twitter covariates provides a granular view of retention correlates.
+Notably, the association with interaction network diversity remains robust ($HR=0.86,p<0.001$), indicating that cross-server engagement predicts retention independent of a user’s prior status on Twitter.
+Among covariates used, the association of metrics extracted from Mastodon remained largely consistent with the first model (see Figure 6).
+Initial followers on the platform ($HR=0.72,p<0.0005$) and number of posts ($HR=0.77,p<0.001$) were still strong factors linked to retention.
+Similarly, server type ($HR=0.79,p<0.001$) and discoverability ($HR=0.86,p<0.001$) remained significant and positively associated with persistence.
+However, unlike the Mastodon-centric model, server structural metrics and Multidisciplinary scholar indicator were no longer significant predictors of attrition in the cross platform model.
+Results of modeling covariates from Twitter revealed nuanced effects of some of the cross-platform user characteristics on persistence.
+A large number of Twitter following ($HR=1.13,p<0.0005$) and followers ($HR=1.11,p<0.0005$) were significant risk factors for becoming inactive on Mastodon.
+In contrast, a higher number of Twitter posts ($HR=0.92,p<0.001$) and older Twitter account age ($HR=0.88,p<0.0005$) were linked to higher retention.
+The results from the cross-platform analysis shows a clear relationship with Twitter history.
+A large Twitter network is associated with higher attrition, while a long history of Twitter activity is linked to persistence showing divergent associations for these factors.
+Users with a well-established and high posting volume on Twitter were more committed to migration but the size of their network may have been too difficult to replicate on Mastodon pulling them back to Twitter.
+This suggests that a successful migration of a community from a well-established platform like Twitter requires not only individual commitment but also a critical mass of connections that can migrate to the new platform.
+Figure 6: Plot showing hazard ratio and 95% CI for each covariate in the cross-platform model. Vertical dashed line at HR of 1.0 represents no effect.
+Contextual Factors and Platform Environment
+The model identifies measurable predictors of sustained use, yet migration and retention dynamics also unfold within a broader contextual environment that is not fully observable from activity traces alone.
+In this section, we examine three such contextual factors: the emergence of competing alternative platforms, public discourse surrounding the academic migration, and Mastodon’s platform affordances.
+These factors help contextualize the conditions under which early adopters navigated migration and sustained participation.
+Alternatives to Mastodon
 In addition to Mastodon, other centralized alternatives to Twitter have emerged in recent years, and many academics explored these venues.
@@
 Specifically, we focused on mentions of Bluesky [^9] https://bsky.app/ and Threads [^10] https://www.threads.net/ using keyword matching as well as user name pattern matching.
-For Mastodon and Bluesky, we also included scholars who use decentralized platform-style usernames as their Twitter usernames (e.g., ¡username¿@¡servername¿).
+For Mastodon and Bluesky, we also included scholars who use decentralized platform-style usernames as their Twitter usernames (e.g., <<username>>@<<servername>>).
 Figure 7 shows an UpSet plot illustrating the number of users who explicitly mention each platform or a combination of platforms.
@@
 This competition further complicates Mastodon’s ability to establish itself as the primary platform for academic discourse.
-Figure 7. Mentions of Platforms in Username and Profile Description.
-In summary, we observe that the majority of scholars are either solely using Twitter or using both Twitter and Mastodon simultaneously.
-Comparing the follower and following counts across these platforms, Mastodon hosts a smaller community.
-Competition from other emerging social media platforms further limits Mastodon’s growth, indicating that scholars are exploring various alternatives rather than committing to Mastodon.
-4.4. Analysis of Tweets on Mastodon Migration
-To further investigate the conversations surrounding Mastodon among academics, we leverage data collected from Twitter reflecting conversations about academics on Mastodon (see Section 3).
-We plot the daily frequency of these tweets and engagement counts over time aggregated by week in Figure 8.
-Discussions about Mastodon in the academic context experienced a noticeable spike during the first two months of the movement.
-After the migration period, although conversations were less frequent, they received more attention (likes), indicating that the audience and connections of academics concerned about Mastodon have largely remained active on Twitter.
-Conversations reflect difficulties and concerns during the transition. For example, Mastodon struggled to accommodate the abrupt influx of users. On the peak day of discussion, November 18, 2022:
-”I have had a Mastodon presence for several months but haven’t used it. I still have a vain hope Twitter will sort itself out, but if not, I am ¡scholar’s Mastodon account¿.”
-”As you know, I’m over on Mastodon too.
-Personally I don’t think that I will leave Twitter until and unless Elon finally and truly kills it, but I know that we will connect over on mastodon if he does.”
+Figure 7: UpSet plot summarizing explicit cross-platform self-identification; intersection bars indicate counts of scholars mentioning Mastodon, Bluesky, and Threads (alone or jointly) in their profiles.
+Public Discourse on Academic Migration
+To examine public discourse surrounding early academic migration, we analyze public posts from Twitter and Mastodon that explicitly reference academic migration, collected using a predefined keyword-based retrieval strategy (see Section Data).
+We plot the daily frequency of these posts and engagement counts over time aggregated by week in Figure 8.
+We observe similar patterns on both platforms: discussions about the migration in the academic context experienced a noticeable spike during the first two months of the movement on both platforms.
+After the migration period, the conversations were less frequent on both platforms, they received more attention (favorites) on Twitter, indicating that the audience and connections of academics concerned about Mastodon have largely remained active on Twitter.
+Qualitative inspection of highly engaged posts during peak periods reveals recurring themes related to uncertainty about platform stability and long-term viability.
+For example, Mastodon struggled to accommodate the abrupt influx of users.
+On the peak day of discussion, November 18, 2022:
+“I have had a Mastodon presence for several months but haven’t used it. I still have a vain hope Twitter will sort itself out, but if not, I am <<scholar’s Mastodon account>>.” –Twitter Post
 However, this initial surge in discussion was short-lived. The frequency of related tweets sharply declined.
-Figure 8. Dynamic frequency of Twitter conversations related to academics migrating to Mastodon and corresponding engagement metrics (connection and like counts), aggregated by week.
+Figure 8: Dynamic frequency of Twitter and Mastodon conversations related to academics migration and corresponding engagement metrics (public interaction and favorite counts), aggregated by week.
 Users expressed frustration over the platform’s lack of centralized theme and purpose, and its fragmented community structure.
-”Vibe-wise, mastodon has *a ton* of different servers already, so there’s more diversity.
-but the *big* servers that most ppl are on are… kinda academic?
-kinda techy?
-very… not fun?
-it feels like linkedin or facebook groups to a lot of people.”
-”My humanities folks have largely migrated to mastodon, but my science and public health feeds are still here on Twitter. More’s the pity.”
-Despite these challenges, ongoing efforts to promote Mastodon indicate a persistent belief in its potential. For instance, in June 2023, one scholar posted,
-”Reasons to join us in the fediverse.
+“My humanities folks have largely migrated to mastodon, but my science and public health feeds are still here on Twitter. More’s the pity.” –Twitter Post
+Despite these challenges, public efforts to encourage a shift toward Mastodon continued throughout 2023. For example, one scholar highlighted its advantages in a June 2023 post on Twitter:
+“Reasons to join us in the fediverse.
 500 character limit.
@@
 Lots more #openaccess resources for #teachers, researchers, science communicators, lifelong learners.
-#STEM #education #openscience #K12 #HigherEd”
-Amongst the discussions, we again observe concurrent mentions of other platforms such as Bluesky and Threads.
+#STEM #education #openscience #K12 #HigherEd” –Twitter Post
+Momentum was also visible within Mastodon itself. A widely circulated post from October 2023, which received 613 reblogs, signaled a clear departure from Twitter:
+Farewell, [bird emoji]!
+After 14 years on Twitter/X, we feel it’s time to part ways.
+Recent developments there do not align with our core values.
+We stand for trustworthy information provision and are committed to open science.
+Follow us here for open science-related updates and news on our services and support for the academic community.[heart emoji] #OpenScience #AcademicLibrary #GoodbyeTwitter #ResearchSupport #EducationSupport –Masotdon Post
+Amongst the discussions on Twitter, we again observe concurrent mentions of other platforms such as Bluesky and Threads.
 These have diluted the migration as users explore various platforms for different purposes.
 For instance,
-”I’m curious re categories of people I follow and where they go.
-IT people seem to have mostly gone to mastodon; librarians ditto; academic leftists stayed here longest, now going to bluesky (prefer commercial sites?); archaeologists have vanished for me completely (threads?).”
-”Hey if anyone’s still here on science Twitter and wondering where a lot of people have gone, they’re trying out bluesky and liking it. This time it’s real (not like mastodon). #academic”
-From a qualitative perspective, these conversations reflect both optimism in Mastodon’s potential and the practical concerns academics face during the migration.
-5. Discussion and Conclusion
+“Hey if anyone’s still here on science Twitter and wondering where a lot of people have gone, they’re trying out bluesky and liking it.
+This time it’s real (not like mastodon).
+#academic” –Twitter Post
+Taken together, these posts illustrate a mix of optimism about Mastodon’s potential and recurring practical concerns users have faced during the migration.
+Mastodon’s Technical Affordances
+Beyond the individual- and community-level factors identified above, structural characteristics of Mastodon, such as its moderation policies and discoverability mechanisms, can be considered additional factors contributing to the limited success of migration.
+In line with the network structure, moderation policy is also decentralized, as described [^11] https://docs.joinmastodon.org/admin/moderation/:
+Moderation in Mastodon is always applied locally, i.e.
+as seen from the particular server.
+An admin or moderator on one server cannot affect a user on another server, they can only affect the local copy on their own server.
+This decentralized moderation stands in sharp contrast to Twitter’s centralized enforcement system.
+While local control empowers individual communities to define and enforce their own standards, it also creates inconsistencies across the network.
+For users accustomed to a uniform, centralized policy environment such as Twitter’s, these variations may generate uncertainty or limit trust in the platform.
+Discoverability presents another technical challenge as shown in the models.
+Unlike Twitter, which supports platform-wide search and algorithmic amplification of content, Mastodon limits search to hashtags and profile information  [^12] https://docs.joinmastodon.org/user/network/#search.
+As a result, newcomers may struggle to discover communities, topics, or influential users, creating barriers to community formation and sustained engagement.
+Taken together, these limiting affordances could help explain why Mastodon, despite its alignment with values of autonomy and community governance, has struggled to attract and retain large numbers of migrating users from Twitter.
+Importantly, these technical affordances are not experienced uniformly across the Fediverse.
+Because each Mastodon server operates with its own administrative history, governance structure, and moderation capacity, users’ experiences may vary substantially depending on the server they join.
+Servers established prior to the migration period may benefit from more mature moderation practices or greater administrative experience, which could influence user satisfaction and retention.
+Discussion and Conclusions
 In this work, we explored the migratory dynamics of Academic Twitter communities to Mastodon.
 Through a series of analyses, we highlight that building a community on a new platform requires more than a coordinated effort and availability of alternatives.
-Results from longitudinal analysis of Mastodon user data showed that the initial surge in sign-ups did not translate into sustained long-term user engagement.
-Most recent user activity on Twitter reveals that many academics in our dataset returned to using Twitter actively.
-Further analysis of engagement metrics reveals that the level of established history, as well as the strong communities established on Twitter, with some over a decade, proved too significant to overcome.
-Without a complete migration of user history, and established community, professionals will naturally gravitate back to familiar platforms where their communities are already established through years.
-Several factors likely contributed to the failure of Mastodon to retain users.
-First, its decentralized structure comes with a steep learning curve compared to Twitter.
-Furthermore, competition from more user-friendly platforms like Bluesky and Threads potentially drew users away from Mastodon.
-Additionally, those who continued to post consistently on Mastodon, did not receive same level of engagement as on Twitter.
-While some users expressed doubts, certain groups of academics stayed committed to Mastodon, making this migration less of a straightforward failure and more of a complex transition.
-One significant limitation of this study is the inconsistency in data collection timing across different platforms.
+Results from longitudinal analysis showed that the initial surge in sign-ups did not translate into sustained long-term user engagement.
+This pattern is consistent with prior studies on Twitter to Mastodon migration in general user groups showing early migration waves often do not translate to stable long term activity with many users maintaining cross platform presence and continue using Twitter as a primary platform for engagement rather than fully switching (Jeong et al.
+2023; He et al.
+2023) and with broader findings that the transition away from Twitter remains incomplete for general audiences (Peña-Fernández et al.
+2025).
+Most recent user activity on Twitter revealed that many academics in our dataset returned to using Twitter.
+Further analysis suggested that the established history and strong communities established on Twitter, with some over a decade, proved too significant to overcome.
+While prior studies observed this general return t Twitter (Jeong et al.
+2023; He et al.
+2023), our cross-platform analysis with Twitter features, suggest that for established scholars, the difficulty of replicating existing networks is correlated with attrition.
+Without a complete migration of user history and established community, professionals and other specialized groups may naturally gravitate back to familiar platforms where their communities are already robust.
+Likewise, several factors were linked to the failure of Mastodon to retain users.
+First, its decentralized structure creates a steeper learning curve compared to Twitter.
+Competition from more user-friendly platforms like Bluesky and Threads perhaps drew users away from Mastodon.
+Beyond direct Twitter substitutes, professional platforms such as LinkedIn continue to serve distinct functions centered around academic communication (Utz and Breuer 2019), which may further diffuse scholars’ attention across platforms rather than concentrating migration on a single alternative.
+Our analysis adds to the understanding of challenges in migration of professional communities by identifying interaction diversity as a robust predictor of retention.
+Additionally, those who continued to post consistently on Mastodon did not enjoy the same level of engagement as on Twitter.
+That being said, select groups of academics stayed committed to Mastodon, making this migration less of a straightforward failure and more of a complex transition.
+Prior work suggests that social influence and a shared sense of identity or belonging can shape collective migration behavior (Cava et al.
+2023).
+In our data, higher persistence among scholars on field or topic specific servers is consistent with the possibility of community specific servers supporting sustained engagement.
+This commitment is likely reinforced by self selection.
+For example, the most motivated early adapters may have actively sought out community specific servers.
+Future efforts that support migrations to decentralized platforms must address these core issues, foster community building by reducing friction, and consider network effects on user engagement.
+While the idea of decentralization of social media platforms holds significant promise, practical challenges persist that require more than just technical solutions.
+The Academic Twitter community’s large scale attempt to migrate from Twitter to Mastodon is significant–a strong professional network with a shared purpose positioned it as a hopeful candidate to succeed with such migration.
+Yet, as encouraging as these efforts and their partial successes are, evidence suggests that the process of transition is extremely challenging.
+Limitations
+One limitation of this study is inconsistency in the timing of data collection across platforms.
 This discrepancy could potentially introduce biases in the analysis, as activity levels of scholars might vary depending on when the data was captured.
 Additionally, incorrect or incomplete social media handles provided by the scholars led to gaps in data that we could not manually resolve due to the scale of the dataset.
-Moreover, the changes to platform access (such as Twitter API rate limitations) restricted our ability to gather sufficient data throughout the study.
-However, we believe our findings from the remaining data is substantial and provides valuable insights for the cross platform analyses.
-Future efforts that support migrations to decentralized platforms must address these core issues, foster community building by reducing friction, and considering network effects to retain user engagement.
-While the idea of decentralization of social media platforms holds a significant promise, practical challenges persist that require more than just technical solutions.
-However, the Academic Twitter community’s large scale attempt to migrate from Twitter demonstrates their adaptability and resilience.
-If any group could successfully make this migration, it would be a community like Academic Twitter, united by a shared purpose and strong networks.
-Yet, as encouraging as these efforts are, there is a long way to go.
+While this subset provides substantial coverage of early adopters and enables examination of cross-platform engagement behaviors within this group, the analysis is descriptive in nature and not intended to support population-level claims.
+Moreover, changes to platform access (e.g., Twitter API rate limitations) restricted our ability to gather complete data throughout the study.
+Similarly, full text keyword search is unavailable on many Mastodon instances unless it is enabled by the server administrator, so our reliance on hashtag-based search, which works without special server configuration, may have resulted in a smaller set of retrievable posts.
+In addition, the keywords used to collect public discourse may not capture all forms of academic self-description, which introduces the possibility of missing some relevant posts.
+However, we believe our findings are nevertheless substantial and provides valuable insights for cross-platform analyses.
+Notably, collected data reflects early groups of academic migrants, which are likely more indicative of collective migration efforts.
+However, we acknowledge the possibility of selection bias, as these patterns may not represent later migrants.
+This focus on early adopters implies that our findings primarily capture the initial, coordinated dynamics of migration rather than the full range of user experiences.
+Furthermore, our analysis treats a user’s server as static which does not account for potential between-server migration within the Fediverse after the server choice at account creation.
+Finally, there may be case-specific factors at both the individual and community levels, including differences in local governance, historical development, or disciplinary norms, that cannot be captured holistically through platform traces alone.
+While such factors fall outside the scope of this study, we believe the observed patterns nevertheless provide a useful empirical account of early academic migration dynamics.
+Ethical Considerations
+All account IDs in the original curated lists were voluntarily added by the users.
+Our data collection was restricted to publicly available interaction and profile metadata on Mastodon and Twitter at the time of collection.
+No private posts, direct messages, or other non-public information were accessed or retained.
+All identifiable information has been removed from the paper and associated research artifacts to further protect user privacy.
+Given the relatively narrow nature of the data collected in this study, we believe the risk of privacy violation and misuse to be minimal.
+Appendix
+Field Discipline Count Total
+Arts and Humanities Archaeology 126 1522
+Art and Design 8
+Biblical Studies 4
+Digital Humanities 208
+Glams (Galleries, Libraries, Archives, Museums, Special Collections) 109
+Literature Studies 142
+Philosophy 220
+Religion 56
+Theology 27
+History 554
+History and philosophy of science 11
+BookHistory 36
+Medieval studies 21
+Business Marketing 4 4
+Education Education 244 244
+Law Law 587 587
+Life Sciences Bioinformatics 13 84
+Biophysics 8
+Palaeogenomics 6
+Plant Science 57
+Medicine and Health Sciences Neuroscience 264 287
+Medical AI 23
+Physical Sciences and Mathematics Chemistry 240 851
+Cheminformatics and Computational Chemistry 11
+Geology 145
+Meteorology 93
+Nuclear fusion 1
+Astrophysics 53
+Atmospheric and air quality 41
+Dendrochronology 3
+Information Security 264
+Social and Behavioral Sciences Anthropology 50 2918
+Communication and Media Studies 265
+Criminology 41
+Genealogy 168
+Geography 282
+Linguistics 108
+Political Science 364
+Sociology 228
+Asian Studies 26
+Iran and middle eastern studies 5
+Psychology 321
+Public Policy 30
+Technology Policy 152
+Science and technology studies 41
+Semantic Web 25
+Qualitative Research 35
+Journalism 777
+Multi-disciplinary Multi-disciplinary 961 1040
+Open Science 79
+Table A1: Number of scholars by disciplines within major academic fields.
+Field One Time(%) Short Term(%) Long Term(%) Persistent(%) Total Percent In-Degree Out-Degree
+Social & Behavioral Sciences 25.0 25.8 24.1 25.1 2918 38.72 98881 91958
+Arts & Humanities 29.0 21.4 26.0 23.6 1522 20.19 38692 39986
+Physical Sciences & Mathematics 22.1 23.4 27.6 26.9 851 11.29 11453 11980
+Life Sciences 20.2 22.6 35.7 21.4 84 1.11 338 648
+Education 32.8 25.0 20.5 21.7 244 3.24 2249 4171
+Law 29.6 28.3 25.4 16.7 587 7.79 21664 27450
+Medicine & Health Sciences 24.7 31.0 26.8 17.4 287 3.81 4292 4923
+Multidisciplinary 18.6 25.1 27.1 29.2 1040 13.80 56371 52818
+Business 25.0 50.0 0.0 25.0 4 0.05 60 66
+Table A2: User activity patterns by academic field (one-time, short-term, long-term, persistent) and field-level statistics (total users, percentage, and network centrality indicators).
+Server Count Percentage In-Degree Out-Degree Description Scope
+mastodon.social 1606 21.31% 147295 128525 The original server operated by the Mastodon gGmbH non-profit. Universal
+mstdn.social 538 7.14% 9456 11947 A general-purpose Mastodon server with a 500 character limit. All languages are welcome. Universal
+mastodon.online 429 5.69% 6600 7172 A newer server operated by the Mastodon gGmbH non-profit. Universal
+fediscience.org 415 5.51% 8220 9628 Fediscience is the social network for scientists. Universal
+mas.to 381 5.06% 5591 6534 Hello! mas.to is a fast, up-to-date and fun Mastodon server. Universal
+sciences.social 218 2.89% 5535 7460 Non-profit, ad-free social media for social scientists. Join thousands of social scientists here and across the fediverse. Field-level
+infosec.exchange 200 2.65% 1001 1362 A Mastodon instance for info/cyber security-minded people. Discipline-level
+scholar.social 193 2.56% 2842 3899 Microblogging for researchers, grad students, librarians, archivists, undergrads, high schoolers, educators, research assistants, profs—anyone involved in learning who engages with others respectfully.
+Universal
+journa.host 165 2.19% 3086 2717 The server for working journalists and news outlets on Mastodon.
+Home to active & retired journalists, media scholars, and a variety of news and journalism adjacent professionals.
+#Newstodon Discipline-level
+zirk.us 153 2.03% 2276 3087 Literature, philosophy, film, music, culture, politics, history, architecture: join the circus of the arts and humanities!
+For readers, writers, academics or anyone wanting to follow the conversation.
+Field-level
+mastodon.world 144 1.91% 1655 2213 Generic Mastodon server for anyone to use. Universal
+toot.community 107 1.42% 805 1202 A worldwide Mastodon instance from The Netherlands. Run by digital enthusiasts, inviting everyone, everywhere to join us in the #fediverse. Universal
+hcommons.social 121 1.61% 2110 2363 hcommons.social is a microblogging network supporting scholars and practitioners across the humanities and around the world. Universal
+aus.social 101 1.34% 1156 1255 Welcome to thundertoot! A Mastodon Instance for the People. Universal
+Table A3: Centrality metrics for Mastodon servers with more than 100 scholars, ordered by count.
+Figure A1: Monthly average total public metrics and statuses count for active users with 10–90% Range.
+Figure A2: Internal interaction networks among servers with more than 100 scholars. Left: boost; Middle: favorite; Right: mention.
+Figure A3: internal follower-followee network for Information Security community (removed isolates for better visualization).
+Figure A4: Connection metrics mapping and differences in metrics between mastodon and Twitter.
+Figure A5: Scatter Plot of Scholar’s Twitter Account Creation Date VS. Engagement Metrics.
+Field/Topic specific servers
+infosec.exchange, journa.host, scholar.social, fediscience.org, sciences.social, zirk.us, hcommons.social, astrodon.social, drosophila.social, fosstodon.org, qoto.org, masto.ai, genomic.social, hci.social, mathstodon.xyz, med-mastodon.com, ecoevo.social, bayes.club, dair-community.social, mapstodon.space, sigmoid.social, universeodon.com, sciencemastodon.com, neuroscience-mastodon.com, datasci.social, meteo.social, mstdn.science, ioc.exchange, wxcloud.social, skew-t.com, eupolicy.social, law.builders, mastodon.lawprofs.org, esq.social, legal.social, publiclaw.space, openbiblio.social, ausglam.space, lingo.lol, archaeo.social, medievalist.masto.host, econtwitter.net, digipres.club, glammr.us, c18.masto.host, religion.masto.host, fedihum.org, bildung.social, book_historia, coastalhistory, fediphilosophy.org, newsie.social, writing.exchange, urbanists.social, mastodon.gamedev.place, peoplemaking.games, photog.social, mastodon.education, educator.social, mastodon.oeru.org
+Table A4: List of field specific servers.
+(a) Primary Model: 30-Day Entropy Window (Concordance: 0.69)
+(b) 14-Day Entropy Window (Concordance: 0.68)
+(c) 60-Day Entropy Window (Concordance: 0.69)
+(d) First 50 Interactions (Concordance: 0.70)
+Figure A6: Mastodon-Centric Model Sensitivity Analysis.
+Forest plots displaying hazard ratios for predictors of inactivity across different specifications of the the federation interaction network metric.
+The protective effect of network diversity and multidisciplinary status remains robust across all time windows.
+Note that Locked Account status becomes a significant risk factor only in the 60-day window ($p<0.05$), suggesting a cumulative effect.
+(a) Primary Model: 30-Day Entropy Window (Concordance: 0.69)
+(b) 14-Day Entropy Window (Concordance: 0.70)
+(c) 60-Day Entropy Window (Concordance: 0.69)
+(d) First 50 Interactions (Concordance: 0.70)
+Figure A7: Cross-Platform Model Sensitivity Analysis.
+Forest plots for the subset of users with Twitter history.
+The plots confirm that Federated Interaction Diversity predicts retention independent of Twitter metadata.
+Unlike the Mastodon-centric model, the Multidisciplinary indicator is not statistically significant in this subset, while Twitter-derived metrics (Followers, Account Age) show consistent divergent associations with retention.